### v2

# Random harmonic maps into spheres

Abstract.
Let $S$ be a punctured Riemann surface with Euler characteristic $\chi(S)<0$. For any unitary representation $\rho:\pi_{1}(S)\to U(N)$, we introduce its renormalized energy and its harmonic representatives, which are equivariant harmonic maps from the universal cover of $S$ to the unit sphere in $\mathbb{C}^{N}$. Our main result is that if a sequence of unitary representations $\rho_{j}$ strongly converges, then their renormalized energies converge to $\frac{\pi}{4}|\chi(S)|$ and the shape of their harmonic representatives converges to a unique rescaled hyperbolic metric. Combining this statement with examples of strongly converging representations provided by random matrix theory, we derive the following applications.
  • If $\pi_{1}(S)$ is a free group, then for a random $\rho:\pi_{1}(S)\to U(N)$, the shape of its harmonic representatives concentrates around a rescaled hyperbolic metric with high probability as $N\to\infty$.
  • For any closed hyperbolic surface, a finite covering admits a harmonic immersion into some Euclidean unit sphere, which is almost isometric after rescaling.
  • There are closed, branched, minimal surfaces $\mathfrak{S}_{j}$ in some Euclidean unit spheres such that $\mathfrak{S}_{j}$ Benjamini-Schramm converges to a rescaled hyperbolic plane as $j\to\infty$, and the Gaussian curvature $K_{j}$ of $\mathfrak{S}_{j}$ satisfies
    $\lim_{j\to\infty}\frac{1}{\operatorname{Area}(\mathfrak{S}_{j})}\int_{\mathfrak{S}_{j}}|K_{j}+8|=0.$
Introduction
A smooth map from a Riemann surface to a Riemannian manifold is harmonic if it is a critical point of the Dirichlet energy functional. Since the seminal paper by J. Eells and J. Sampson [ES64], tremendous effort has been devoted to equivariant harmonic maps into nonpositively curved spaces coming from representations of surface groups: given a Riemann surface $S$ and a representation of its fundamental group into the isometry group of a Riemannian manifold $X$ with nonpositive curvature,
$\rho:\pi_{1}(S)\to\operatorname{Isom}(X),$
one can often construct by energy minimization a harmonic map from the universal cover of $S$ to $X$ which is $\pi_{1}(S)$-equivariant with respect to $\rho$. These equivariant harmonic maps are rigid, for instance they are unique up to natural equivalence [Har67]. This type of rigidity plays a fundamental role in the non-abelian Hodge correspondence [GP09] [Li19][Tho24], in Teichmüller theory [DW07] [JY09] [Lab19], and has been a driving force in the development of geometric analysis [SY97][Jos08].
What happens when the target space $X$ is not nonpositively curved? Consider equivariant harmonic maps into Euclidean spheres. While rigidity in the traditional sense does not hold, we find a new kind of rigidity which is more probabilistic in nature: when the dimensions of the spheres grow to infinity, the shapes of equivariant harmonic maps tend to concentrate around a unique hyperbolic geometry.
To formalize this statement, we combine harmonic map theory with the topic of unitary representations of surface groups. This provides a playground for exploring the effect of randomness and high dimensionality on variational objects such as harmonic maps and minimal surfaces. In particular, we establish a concrete connection between random matrix theory and harmonic map theory, which is the content of the main theorem.
0.1. Renormalized energy and equivariant harmonic maps
Consider a punctured Riemann surface $S$, namely a closed oriented surface minus finitely many points and endowed with a conformal structure. Suppose that $S$ has negative Euler characteristic $\chi(S)<0$. For $N\geq 1$, consider a unitary representation of its fundamental group:
$\rho:\pi_{1}(S)\to U(N).$
The representation $\rho$ induces an isometric action of $\pi_{1}(S)$ on the unit sphere $(\mathbb{S}^{2N-1},g_{\mathrm{Eucl}})$ in $\mathbb{C}^{N}$ with its standard Euclidean metric $g_{\mathrm{Eucl}}$.
One of the simplest geometric invariants for the pair $(S,\rho)$ is its energy:
[…]
where $\tilde{S}$ is the universal cover of $S$, $\pi_{1}(S)$ acts by deck transformations on $\tilde{S}$, $\mathbf{D}_{S}\subset\tilde{S}$ is any Borel fundamental domain of this action, and the $L^{2}$-norm of the differential of $\varphi$ is computed with respect to $g_{\mathrm{Eucl}}$ on $\mathbb{S}^{2N-1}$ and any Riemannian metric on $S$ compatible with it conformal structure.
When $S$ has punctures, $\operatorname{E}(S,\rho)$ is typically infinite. By applying a standard renormalization procedure that originated in the work of F. Bethuel, H. Brezis and F. Hélein [BBH+94], we define our main invariant, the renormalized energy of $(S,\rho)$:
[…]
It is a finite number and is obtained by minimizing the energy after subtracting off the diverging logarithmic energy contribution near the punctures (see Subsection 1.1). When $\operatorname{E}(S,\rho)$ is finite, we have $\operatorname{E}(S,\rho)=\operatorname{E_{\mathrm{ren}}}(S,\rho)$.
The main geometric property of the renormalized energy $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ is that it is always “achieved” by an optimal harmonic map
$\psi:\tilde{S}\to(\mathbb{S}^{2N-1},g_{\mathrm{Eucl}})$
which is $\pi_{1}(S)$-equivariant with respect to $\rho$ and equivariantly energy-minimizing, see Theorem 1.7. Any such $\psi$ will be called a harmonic representative of $(S,\rho)$. Its intrinsic shape is encoded by the pullback metric $\psi^{*}g_{\mathrm{Eucl}}$ on $\tilde{S}$, which descends by equivariance of $\psi$ to a Riemannian metric on $S$, still called $\psi^{*}g_{\mathrm{Eucl}}$.
How does the renormalized energy behave? What is the geometry of the harmonic representatives? We will answer these questions in the large $N$ limit, when the representation $\rho$ is chosen at random.
0.2. Applications
Before stating our main result, here are some unexpected corollaries.
Geometric concentration: The unique conformal, finite-area, complete hyperbolic metric on $S$ is denoted by $g_{\mathrm{hyp}}$. Suppose that $\pi_{1}(S)$ is isomorphic to a nonabelian free group $F_{k}$ of rank $k\geq 2$. Since $\chi(S)<0$, this occurs if and only if $S$ has at least one puncture. Unitary representations of $\pi_{1}(S)$ into $U(N)$ are then in one-to-one correspondence with $k$-tuple $(u_{1},...,u_{k})\in U(N)^{k}$, and the Haar measure on $U(N)^{k}$ provides a natural notion of random unitary representation of $\pi_{1}(S)$, see Subsection 5.1. Our first application is a geometric concentration result for the shape of harmonic representatives:
Theorem 0.1.
Let $\epsilon>0$ and let $K\subset S$ be a compact domain. If $N$ is large enough, then for a random unitary representation $\tau:\pi_{1}(S)\to U(N)$ and any harmonic representative $\psi$ of $(S,\tau)$, with probability at least $1-\epsilon$,
$|\operatorname{E_{\mathrm{ren}}}(S,\tau)-\frac{\pi}{4}(k-1)|<\epsilon\quad\text{and}\quad\|\psi^{*}g_{\mathrm{Eucl}}-\frac{1}{8}g_{\mathrm{hyp}}\|_{C^{2}(K)}<\epsilon.$
Here, the $C^{2}$-norm is computed with respect to $g_{\mathrm{hyp}}$. Thus, $\operatorname{E_{\mathrm{ren}}}(S,\tau)$ tends to concentrate around a value independent of the Riemann surface $S$, once the rank of $\pi_{1}(S)$ is fixed, while the pullback metric by harmonic representatives tends to concentrate around a rescaled hyperbolic metric which determines the Riemann surface $S$.
Special immersions of surfaces into Euclidean spaces: Our next corollary concerns an old theme in differential geometry:
What is the best way to immerse a closed surface inside a Euclidean space?
Consider a closed hyperbolic surface $(\Sigma,g_{\mathrm{hyp}})$. Nash’s embedding theorem provides a massive collection of isometric immersions of $(\Sigma,g_{\mathrm{hyp}})$ into Euclidean spaces [And02]. On the other hand, it was later shown that such an isometric immersion of $(\Sigma,g_{\mathrm{hyp}})$ never satisfies other natural conditions: it cannot be a harmonic map into a Euclidean sphere [Bry85, Theorem 2.3], or have parallel mean curvature [CL72, Theorem 2] [CC73, Classification Theorem] [Yau74, Theorem 4] or have parallel normalized mean curvature [Che80, Theorem 2]. Since then, determining the boundary between these two regimes of extreme flexibility and rigidity had been open. Our second application provides an essentially optimal answer:
Theorem 0.2.
Let $\epsilon>0$. For any closed hyperbolic surface $(\Sigma,g_{\mathrm{hyp}})$, there exists a finite degree covering $(\Sigma^{\prime},g_{\mathrm{hyp}})$ and a harmonic immersion into a Euclidean unit sphere
$\psi:(\Sigma^{\prime},g_{\mathrm{hyp}})\to(\mathbb{S}^{n_{\epsilon}},g_{\mathrm{Eucl}})$
such that $\|\psi^{*}g_{\mathrm{Eucl}}-\frac{1}{8}g_{\mathrm{hyp}}\|_{C^{2}(\Sigma^{\prime})}<\epsilon$.
Here the $C^{2}$-norm is computed with respect to $g_{\mathrm{hyp}}$. We speculate on the meaning of the factor $\frac{1}{8}$ in Remark 5.3. These harmonic immersions are obtained using random constructions, and as of now there is no known construction of such maps based on more conventional, deterministic methods. To our knowledge, this is the first application of the probabilistic method and random matrix theory to special immersions of surfaces.
Almost hyperbolic minimal surfaces in spheres: For our third application, recall that a closed, branched, minimal surface is the image of a closed surface by a harmonic map which is weakly conformal. The study of minimal surfaces in spheres has a long history. Minimal round spheres and minimal flat tori in Euclidean spheres exist, and are completely classified [Cal67] [Ken76][Bry85, Proposition 3.3]. On the other hand, the obstruction of R. Bryant [Bry85, Theorem 2.3] states that minimal surfaces in Euclidean spheres can never be hyperbolic surfaces [Bry85, Theorem 2.3]. This obstruction is one of the few existing results on the relation between minimal surfaces in spheres and negative Gaussian curvature. Curiously, the following problem from S.-T. Yau’s 1982 list [Yau82, Problem Section, Problem 101] is still open:
Is there a closed, minimal surface $\mathfrak{S}$ in a Euclidean sphere with negative Gaussian curvature $K_{\mathfrak{S}}<0$?
As a corollary of the main theorem, we show the existence of new minimal surfaces in high dimensional spheres, which are almost hyperbolic on average:
Theorem 0.3.
There exist a sequence of closed, branched, minimal surfaces $\mathfrak{S}_{j}$ in Euclidean unit spheres $(\mathbb{S}^{n_{j}},g_{\mathrm{Eucl}})$ such that
$\lim_{j\to\infty}\frac{1}{\operatorname{Area}(\mathfrak{S}_{j})}\int_{\mathfrak{S}_{j}}|K_{\mathfrak{S}_{j}}+8|=0.$
Moreover, $\mathfrak{S}_{j}$ Benjamini-Schramm converges to $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$ as $j\to\infty$.
In this statement, $K_{\mathfrak{S}_{j}}$ is the Gaussian curvature of $\mathfrak{S}_{j}$ with respect to the induced metric, $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$ is the rescaled hyperbolic plane with Gaussian curvature $-8$, and Benjamini-Schramm convergence captures the asymptotic geometry of $\mathfrak{S}_{j}$ around most points (it will be defined in Subsection 5.3). Although Theorem 0.3 is relevant to [Yau82, Problem Section, Problem 101], its main point lies elsewhere: the qualitative properties of $\mathfrak{S}_{j}$ are drastically different from the familiar examples of minimal surfaces in $\mathbb{S}^{3}$ or $\mathbb{S}^{4}$ [Law70] [Bry82a] [KPS88] [KY10][CS15]. They support the heuristic that negative curvature should be “typical” for minimal surfaces in higher-dimensional Riemannian manifolds.
0.3. Main result: from strong convergence to harmonic maps
Let $S$ be a punctured Riemann surface of genus $\mathbf{g}$ and with $\mathbf{n}\geq 0$ punctures. Suppose that $S$ has negative Euler characteristic $\chi(S)=2-2\mathbf{g}-\mathbf{n}<0$. Its unique conformal, finite-area, complete hyperbolic metric is denoted by $g_{\mathrm{hyp}}$.
Let us state the definition of strong convergence, an important property coming from free probability and random matrix theory. A sequence of unitary representations $\rho_{j}:\pi_{1}(S)\to U(N_{j})$ strongly converges when for any $z\in\mathbb{C}[\pi_{1}(S)]$,
$\lim_{j\to\infty}\|\rho_{j}(z)\|=\|\lambda_{\pi_{1}(S)}(z)\|$
[…]
Theorem 0.4 (Main theorem).
Let $\rho_{j}:\pi_{1}(S)\to U(N_{j})$ be a sequence of unitary representations and, for each $j\geq 1$, let $\psi_{j}$ be a harmonic representative of $(S,\rho_{j})$. If $\rho_{j}$ strongly converges, then
$\lim_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})=\frac{\pi}{4}|\chi(S)|$
and $\psi_{j}^{*}g_{\mathrm{Eucl}}$ converges to $\frac{1}{8}g_{\mathrm{hyp}}$ in the $C^{\infty}$-topology on compact subsets of $S$ as $j\to\infty$.
This theorem connects an analytic condition on norms of linear operators to the geometric behavior of harmonic maps: under this condition, the rescaled hyperbolic metric emerges as the unique limit of pullback metrics [^1] Building on methods of this paper, together with Riccardo Caniato and Xingzhe Li, we show that under the assumptions of Theorem 0.4, not only is the limit pullback metric unique, but the extrinsic limit shape of the harmonic maps is also unique.. I suspect that, under mild assumptions, the conclusion in this theorem should actually be equivalent to strong convergence. The assumption of Theorem 0.4 is often satisfied thanks to the “universality” of strong convergence: a plethora of models of sequences of random unitary representations all strongly converge. There has been a flurry of activity in random matrix theory on this subject in recent years, starting with the fundamental work of U. Haagerup and S. Thorbjørnsen [HT05]. Remarkably, there is no known deterministic construction of strongly converging sequences of finite dimensional unitary representations. Here are a few results from this large body of work (see Subsection 6.3 in the Appendix and references in [Mag24]): based on [HT05], B. Collins and C. Male showed that sequences of unitary representations of free groups built from random Haar unitaries strongly converge [CM14]; in an important paper, C. Bordenave and B. Collins established the strong convergence of sequences of unitary representations of free groups constructed from random permutations [BC19]; more recently, C.F. Chen, J. Garza-Vargas, J. Tropp and R. van Handel discovered that random “stable representations” of free groups strongly converge, a result which recovers previous results on free groups with new proof methods [CGVTvH24]; L. Louder and M. Magee constructed sequences of unitary representations of surface groups which strongly converge [LMH25].
0.4. Related work and comments
In the genealogy of harmonic map theory, our results belong to the broad theme of representations of surface groups into Lie groups and equivariant harmonic maps into symmetric spaces. The literature on this subject is enormous; we refer the reader to [SY97, Chapters III and XIII] [DW07] [JY09] [Jos08, Chapter 9] [GP09] [Li19] [Tho24] and the references therein. Essentially all previous results focused on nonpositively curved target spaces, where the rigidity of equivariant harmonic maps is leveraged to study representations and their moduli spaces, in a fixed dimension. In contrast, what we establish for representations into unitary groups is a “probabilistic rigidity” statement in the regime where the dimension tends to infinity. This eventually hinges on a rigidity result for certain equivariant minimal surfaces in infinite dimensional spheres, see Section 2. Are there analogous phenomena for representations into general Lie groups?
Unitary representations of surface groups is a classical topic on its own [Lab13]. When the surface has punctures, the fundamental group is free. Its unitary representations are then in one-to-one correspondence with tuples of unitary matrices, which have been intensively studied in random matrix theory and free probability, see [Spe17] and the references in the Appendix. When the surface has no punctures, the unitary representations of its fundamental group are related to stable holomorphic vector bundles by M. Narasimhan and C. Seshadri [NS65]. Their moduli space carries a natural symplectic volume form by M. Atiyah and R. Bott [AB83], and W. Goldman [Gol84], whose volume was computed by E. Witten [Wit91] and A.N. Sengupta [Sen03]. In [Mag21] [Mag22], M. Magee computed the expansion of Wilson loops on this moduli space.
This article is part of a series [Son23b] [Son23a] [Son24] [Son25a] which initiates a study of orthogonal and unitary representations of groups using geometric concepts such as area, energy, minimal surfaces and harmonic maps. The central geometric objects are triples $(\Gamma,\rho,\Sigma)$ where $\Gamma$ is a group, $\rho:\Gamma\to\operatorname{End}(V)$ is an orthogonal representation of $\Gamma$, and $\Sigma$ is a $\rho(\Gamma)$-invariant $m$-dimensional minimal surface in the unit sphere of $V$. When do such triples exist? When are they rigid? The variational problem for such triples is called the spherical Plateau problem. The harmonic maps we consider in this paper are closely connected, since they are solutions of a “spherical Dirichlet problem”, see discussion in Subsection 1.3.
The renormalized energy introduced in this paper is closely connected to several invariants in analysis, representation theory, geometry and topology (see [Son25a] for expanded explanations and related questions):
Bethuel-Brezis-Hélein’s renormalized energy: our renormalized energy is a direct descendant of the renormalized energy introduced by F. Bethuel, H. Brezis and F. Hélein [BBH+94] for harmonic maps from punctured planar domains to the circle. There are numerous generalizations relevant to us, including the generalization to Riemannian targets by A. Monteil, R. Rodiac, J. Van Schaftingen [MRVS22] and the “modified energy” considered by G. Daskalopoulos, C. Mese in [DM23].
Spectral gaps: the renormalized energy can be viewed as a nonlinear analogue of the first Laplace eigenvalue for sections of flat $\mathbb{C}^{N}$-bundles over surfaces [Zar22] [Hid23], and the corresponding harmonic maps into spheres are then nonlinear analogues of first Laplace eigensections.
Kazhdan constants: those invariants [BdLHV08, Remark 1.1.4] measure the distance between a unitary representation $\rho_{\Gamma}$ of a discrete group $\Gamma$ and the trivial representation in the Fell topology. The renormalized energy could be used to define a notion of 2-dimensional Kazhdan constants.
Energy for maps into nonpositively curved spaces: for nonpositively curved spaces, notions of energy or volume of representations analogous to the renormalized energy have been extensively studied [Tol89][Dun99] [Fra04] [BCG07] [Lab17] [PS17]… Harmonic maps with infinite energy appear in [Loh90] [Wol91] [JZ97] [DM23] [Sag23]
Besson-Courtois-Gallot’s spherical volume: one of our original motivations for defining the renormalized energy is the spherical volume of G. Besson, G. Courtois and S. Gallot. It is a topological invariant for any closed manifold $M$, defined by minimizing the “area” of certain maps equivariant with respect to the regular representation of $\pi_{1}(M)$ [BCG91, Section 3].
Minimal surfaces are a special case of harmonic maps. To our knowledge, there are roughly three general methods to construct large genus minimal surfaces in high dimensional spheres: area-minimization under symmetry assumptions [SY79] [SU82], twistor map methods [Bry82a] [Bry82b] [Han96] which produce superminimal surfaces, and $\lambda_{1}$-maximization methods [Nad96] [ESI00] [Pet14] (see also [NS15] [MS19]). The random minimal surfaces of Theorem 0.3 are obtained via the first method. The second method is the most explicit and thus gives more information on the Gaussian curvature of the minimal surfaces. The third method yields minimal surfaces which might potentially [^2] I learnt this possibility from Misha Karpukhin. It seems supported by an estimate of A. Ros [Ros22, Subsection 2.4]. become almost hyperbolic when the genus gets large, like the minimal surfaces in Theorem 0.3. Concerning [Yau82, Problem Section, Problem 101], closed negatively curved minimal surfaces may well already exist in the round 4-sphere $\mathbb{S}^{4}$. Such surfaces do not exist in $\mathbb{S}^{3}$ by [Law70] and they cannot be superminimal surfaces in $\mathbb{S}^{4}$ [Bry82a] by [Bry24].
Our work is closely connected to spectral geometry. Given a Riemann surface $S$ and a representation $\rho:\pi_{1}(S)\to U(N)$, there is a corresponding twisted flat $\mathbb{C}^{N}$-bundle $\mathcal{B}$ over $S$. Equivariant maps from $\tilde{S}$ to $\mathbb{S}^{2N-1}$ descend to sections of $\mathcal{B}$ with constant norm $1$. Thus, the renormalized energy $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ and harmonic representatives are nonlinear analogues of the first Laplace eigenvalue for sections of $\mathcal{B}$ and first Laplace eigensections of $\mathcal{B}$, respectively. One may connect the two sides via Ginzburg-Landau theory, which involves Dirichlet-type energies $E_{\epsilon}$ depending on a parameter $\varepsilon\in(0,\infty)$ (see [BBH+94, Introduction, Equation (1)]). In the $\varepsilon\to 0$ limit, critical points of $E_{\epsilon}$ correspond to harmonic representatives, while in the $\varepsilon\to\infty$ limit, they correspond to Laplace eigensections. In view of Theorem 0.4, what is the limit shape of first Laplace eigensections, for strongly convergent representations? As for the $\lambda_{1}$-maximization methodology [Nad96] [ESI00] [Pet14] [NS15] [MS19], what is the limit shape of metrics on $S$ maximizing the first Laplace eigenvalue, for strongly convergent representations?
In a recent article [HM23], W. Hide and M. Magee settled an old open problem by showing that there are closed hyperbolic surfaces with arbitrarily large genus and with $\lambda_{1}$ arbitrarily close to $\frac{1}{4}$, by studying random coverings of hyperbolic surfaces. Their proof relies on resolvent methods and the strong convergence result of C. Bordenave and B. Collins [BC19], who treated an analogous problem for the spectral gaps of graphs. More recent results [Zar22] [Hid23], based on the method of [HM23], imply that the first eigenvalue of the Laplace operator acting on sections of random flat bundles over a hyperbolic surface $(S,g_{\mathrm{hyp}})$ is at least $\frac{1}{4}-\varepsilon$ with high probability in large dimensions. Those techniques provide [^3] I thank Michael Magee for explaining this to me. sharp lower bounds for the renormalized energy $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ for certain representations $\rho$ in large dimensions, see Remark 4.5. On the other hand, the proof of Theorem 0.4, especially the identification of the limit pullback metric, is based on arguments completely different from [HM23].
The conclusions of Theorems 0.2 and 0.3 are reminiscent of several results about holomorphic curves (which are a special kind of harmonic maps and minimal surfaces) in complex geometry. First, the analogue of [Yau82, Problem Section, Problem 101] for holomorphic curves in complex projective spaces is completely understood thanks to Kodaira embeddings [Tia90][Moh22] and various obstructions [Per] [Hul00]: there are holomorphic curves in complex projective spaces whose induced metric is close to a hyperbolic metric after rescaling. In a similar vein, the period map for Riemann surfaces yields holomorphic curves inside complex tori whose induced metric is close to a hyperbolic metric after rescaling [Kaz70] [Rho93] [McM13]. All those constructions are explicit, in contrast to Theorem 0.3. Is there an explicit construction of almost hyperbolic minimal surfaces in spheres?
Some earlier settings involved various different notions of “random” minimal surfaces. The study of zeroes of random complex polynomials and random holomorphic sections, a special type of minimal submanifolds, is surveyed in [BCHM18] [SZ23]. Minimal surfaces representing surface subgroups in hyperbolic manifolds and other locally symmetric spaces of nonpositive curvature, have been studied using dynamical and topological methods, see [KM12] [Ham15][CMN22] [LN21][KMS23]. Equidistribution and scarring results from min-max minimal surfaces [MNS19] [GG19] [SZ21] [Li23] shed light on the average behavior of minimal surfaces. See also the survey [Son25b].
0.5. Outline of proof for the main theorem
Roughly speaking, given two unitary representations $\tau_{1}$, $\tau_{2}$, we say that $\tau_{1}$ is weakly contained in $\tau_{2}$ if the linear action of $\tau_{1}$ can be arbitrarily well approximated or “imitated” by the linear action of $\tau_{2}$. Weak equivalence between $\tau_{1}$ and $\tau_{2}$ means that one is weakly contained in the other and vice versa, see Subsection 6.1 in the Appendix for definitions. Under the assumptions of the main theorem, there are $\pi_{1}(S)$-equivariant harmonic maps $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ which achieve the renormalized energy $\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})$.
Step 1 (Upper bound for the renormalized energy): The strong convergence of $\rho_{j}$ implies that the direct sum $\bigoplus_{j\geq 1}\rho_{j}$ weakly contains the regular representation $\lambda_{\pi_{1}(S)}:\pi_{1}(S)\to\ell^{2}(\pi_{1}(S))$. For any $\varepsilon>0$, there is an explicit map $\mathscr{P}_{\epsilon}$ from $\tilde{S}$ to the unit sphere of $\ell^{2}(\pi_{1}(S))$, equivariant with respect to $\lambda_{\pi_{1}(S)}$, with energy $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)+\varepsilon$ on a fundamental domain in $\tilde{S}$. By an approximation argument, as $j\to\infty$, we construct maps $\varphi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ equivariant with respect to $\rho_{j}$, with renormalized energy at most $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)+2\varepsilon$. In particular, $\limsup_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})\leq\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$.
Step 2 (Constructing a limit map): The harmonic maps $\psi_{j}$ can be shown to have their standard energy eventually upper bounded by $\limsup_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})+\varepsilon$ for any $\varepsilon>0$, on each fixed compact subset of a fundamental domain of $\tilde{S}$. By the upper bound in Step 1 and standard harmonic map theory, the harmonic maps subsequentially converge on compact subsets to a limit harmonic map $\psi_{\infty}:\tilde{S}\to\mathbb{S}^{\infty}$ inside some Hilbert space. By lower semicontinuity of the energy, the map $\psi_{\infty}$ has energy at most $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$ on a fundamental domain.
Step 3 (Identifying the limit representation): From the limit map $\psi_{\infty}$, we can construct a limit unitary representation $\rho^{\prime}$ with respect to which $\psi_{\infty}$ is equivariant. By using the strong convergence of $\rho_{j}$ a second time (in a much more crucial way), we argue that $\rho^{\prime}$ is in fact weakly equivalent to the regular representation $\lambda_{\pi_{1}(S)}$.
Step 4 (Lower bound for the renormalized energy): We show that any map $\tilde{S}\to\mathbb{S}_{\ell^{2}(\pi_{1}(S))}$ equivariant with respect to $\lambda_{\pi_{1}(S)}$ has energy at least $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$ on a fundamental domain of $\tilde{S}$, and that an analogous property holds for $\rho^{\prime}$ by approximation. We deduce from the previous step that $\psi_{\infty}$ has energy exactly $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$ on a fundamental domain, and that it is actually a branched minimal immersion. It also implies that $\lim_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$ as wanted.
Step 5 (Uniqueness of the limit pullback metric): The last step is to show that any branched minimal immersion $\tilde{S}\to\mathbb{S}^{\infty}$ equivariant with respect to a representation weakly equivalent to $\lambda_{\pi_{1}(S)}$ has to have its pullback metric equal to $\frac{1}{8}g_{\mathrm{hyp}}$. This is proved using a new interpolation argument which takes advantage of the infinite dimensionality of our setting, in addition to the representation theory of $\mathrm{PSL}_{2}(\mathbb{R})$. It finishes the proof since $\psi_{j}$ subsequentially converges on compact subsets to $\psi_{\infty}$.
The key conceptual steps are the use of strong convergence in Step 3, and the uniqueness statement in Step 5. The way we ended up writing our proof is a bit different from that outline; $\lambda_{\pi_{1}(S)}$ will in particular be replaced by another weakly equivalent, explicit, “boundary” unitary representation $\underline{\rho}_{B}$. Once the main theorem is proved, the applications (Theorems 0.1, 0.2, 0.3) follow by using well-chosen pairs $(S,\rho_{j})$.
0.6. Organization of the paper
Section 1: Definition of the energy, renormalized energy and area; construction of equivariant harmonic maps; study of the special case of the regular representation.
Section 2: A general uniqueness result for equivariant minimal surfaces; application to the regular representation.
Section 3: Relation between weak containment and approximation of equivariant maps; constructions of extensions of maps with controlled renormalized energy; application to upper bounds for the renormalized energy.
Section 4: Proof of the main theorem: convergence of the renormalized energy and the pullback metric.
Section 5: Applications of the main theorem to different choices of surfaces and representations.
Section 6: Appendix: definitions of weak containment, weak equivalence and strong convergence; description of the three examples of strongly converging representations which we use.
Acknowlegments
I am grateful to Michael Magee, Ramon van Handel, Benoît Collins and Jorge Garza-Vargas for answering my numerous questions about strong convergence of unitary representations; Robert Bryant for a detailed explanation of his work on minimal surfaces in spheres; Misha Karpukhin and Daniel Stern for interesting exchanges about eigenvalue optimization problems. I would also like to thank Nicos Kapouleas, Xin Zhou, Yuchin Sun and Christine Breiner for helpful discussions about minimal surfaces and harmonic maps; Peter Sarnak, Mario Micallef, Riccardo Caniato and Xingzhe Li for additional conversations.
A.S. was partially supported by NSF grant DMS-2104254. This research was partially conducted during the period A.S. served as a Clay Research Fellow.
1. Renormalized energy and harmonic maps
1.1. Renormalized energy
Let $S$ be a punctured Riemann surface: equivalently, $S$ is a closed oriented surface minus finitely many points called punctures, endowed with a conformal structure. We will almost always assume that $S$ is of general type: it means that its Euler characteristic is negative, or equivalently that it admits a (unique) conformal, complete, finite area, hyperbolic metric called $g_{\mathrm{hyp}}$. Let $\pi_{1}(S)$ be the fundamental group of $S$, let $\tilde{S}$ be its universal cover which is endowed with the lift of $g$, and on which $\pi_{1}(S)$ acts by deck transformations properly and free. Let $\mathbf{D}_{S}\subset\tilde{S}$ be a Borel fundamental domain, which we can choose so that the boundary is piecewise smooth.
Let $V$ be a complex Hilbert space. In this paper, the unit sphere of $V$ will always be called $\mathbb{S}_{V}$. Let
$\rho:\pi_{1}(S)\to\operatorname{End}(V)$
be a unitary representation. It induces an isometric action of $\pi_{1}(S)$ on $\mathbb{S}_{V}$, which is endowed with its standard round Riemannian metric $g_{V}$.
Consider the space of maps
(1) $\mathscr{H}_{S,\rho}:=\{\text{$\pi_{1}(S)$-equivariant smooth maps from $\tilde{S}$ to $\mathbb{S}_{V}$}\}.$
It is an exercise in topology that this set is nonempty. Recall that if $u$ is a smooth map from a Riemann surface $S^{\prime}$ to a Riemannian manifold $(M,g_{M})$, the energy of $u$ restricted to a subset $D\subset S^{\prime}$ is defined as
[…]
where the $L^{2}$-norm of the norm of the differential $|du(x)|$ is computed with respect to any Riemannian metric $g_{S^{\prime}}$ on $S^{\prime}$ compatible with its conformal structure, and the metric $g_{M}$. This is well-defined by conformally invariance of the energy (it remains unchanged after replacing $g_{S^{\prime}}$ by another conformal metric). We drop the subscript $D$ if $D={S^{\prime}}$.
We introduce the spherical energy, or “energy” for short:
Definition 1.1.
The spherical energy of $(S,\rho)$ is defined as
$\operatorname{E}(S,\rho):=\inf\{\operatorname{E}(u|_{\mathbf{D}_{S}});\quad u\in\mathscr{H}_{S,\rho}\}\in[0,\infty].$
When $S$ has at least one puncture, the energy defined above can be infinite in general. In order to get a finite number, we apply the renormalization procedure introduced by Bethuel-Brezis-Hélein [BBH+94] and generalized by Monteil-Rodiac-Van Schaftingen [MRVS22], Daskalopoulos-Mese [DM23]. This procedure depends on a choice of conformal parameterization for a punctured neighborhood of each puncture by a punctured flat unit disk. There are quite a few interesting choices for such parameterizations, but in this paper we will fix the following distinguished choice [^4] Our main theorem does not in fact depend on the choice of parameterization., used in [Wol07] for instance. To define this parameterization, let us recall that in a complete, finite-area, hyperbolic surface, every cusp contains a unique embedded unit area cusp bounded by a curve with constant geodesic curvature, and these unit area cusps are all disjoint. Given $p\in\mathrm{Punc}_{S}$, let $D(p,1)$ be the open region of $S$ surrounding $p$ corresponding to the unit area cusp in the cusp of $p$ when $S$ is endowed with its unique conformal complete, finite area hyperbolic metric $g_{\mathrm{hyp}}$. There is a conformal diffeomorphism unique up to rotations,
$P_{p}:D(p,1)\to\mathbb{D}^{*},$
where $\mathbb{D}^{*}$ is the Euclidean unit disk in $\mathbb{R}^{2}$ minus the origin $O$. For $r\in(0,1]$, let
(2) $D(p,r):=\{q\in D(p,1);\quad\operatorname{dist}_{\mathrm{Eucl}}(P_{p}(q),O)<r\}.$
Denote by $\tilde{D}(p,r)$ the preimage of $D(p,r)$ under the natural projection map $\tilde{S}\to S$.
Given a representation $\rho$ as above and a puncture $p\in\mathrm{Punc}_{S}$, oriented embedded small loops around the puncture determine a conjugacy class $\mathcal{J}_{p}\subset\pi_{1}(S)$. Set
(3) $\operatorname{\boldsymbol{\lambda}}_{\rho}(p):=\inf\{\operatorname{dist}_{\mathbb{S}_{V}}\big{(}x,\rho(J_{p})x\big{)};\quad x\in\mathbb{S}_{V}\}$
where $J_{p}$ is any element in $\mathcal{J}_{p}$ (the definition does not depend on this choice), $\rho(J_{p})x$ denotes the image of $x$ by $\rho(J_{p})$ and $\operatorname{dist}_{\mathbb{S}_{V}}$ is the standard Riemannian distance on $\mathbb{S}_{V}$.
For $u\in\mathscr{H}_{S,\rho}$, its renormalized energy is
$\operatorname{E_{\mathrm{ren}}}(u|_{\mathbf{D}_{S}}):=\liminf_{r\to 0}\big{[}\operatorname{E}(u|_{\mathbf{D}_{S}\setminus\bigcup_{p\in\mathrm{Punc}_{S}}\tilde{D}(p,r)})-\sum_{p\in\mathrm{Punc}_{S}}\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p)^{2}}{4\pi}\log\frac{1}{r}\big{]}.$
Note that if $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ for each $p\in\mathrm{Punc}_{S}$, then $\operatorname{E_{\mathrm{ren}}}(u|_{\mathbf{D}_{S}})=\operatorname{E}(u|_{\mathbf{D}_{S}})$. We can now define the renormalized spherical energy, or “renormalized energy” for short:
Definition 1.2.
The renormalized spherical energy of $(S,\rho)$ is defined as
$\operatorname{E_{\mathrm{ren}}}(S,\rho):=\inf\{\operatorname{E_{\mathrm{ren}}}(u|_{\mathbf{D}_{S}});\quad u\in\mathscr{H}_{S,\rho}\}.$
The renormalized energy is always finite when $\dim_{\mathbb{R}}V<\infty$, see [MRVS22, DM23], or Corollary 3.3 (2). If $S$ is closed (i.e. has no punctures), then $\operatorname{E}(S,\rho)=\operatorname{E_{\mathrm{ren}}}(S,\rho)$.
Remark 1.3 (Invariance).
Both $\operatorname{E}(S,\rho)$ and $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ remain invariant when we conjugate $\rho$ by a unitary operator. The choice of distinguished conformal parameterization for punctured neighborhoods of the punctures is invariant by conformal diffeomorphism. Hence, $\operatorname{E}(S,\rho)$ and $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ induce functions on the product of Teichmüller spaces of surfaces and the moduli spaces of representations of their fundamental group.
Lemma 1.4.
If $\operatorname{E}(S,\rho)$ is finite then $\operatorname{E}(S,\rho)=\operatorname{E_{\mathrm{ren}}}(S,\rho)$.
Proof.
The condition $\operatorname{E}(S,\rho)<\infty$ implies that there is a smooth map $u\in\mathscr{H}_{S,\rho}$ which has finite energy on the fundamental domain $\mathbf{D}_{S}\subset\tilde{S}$. For each puncture $p\in\operatorname{Punc}_{S}$, we use the previous notation $\tilde{D}(p,1)$, the domain $\mathbf{D}_{S}$ can be chosen so that the closure of $\mathbf{D}_{S}\cap\tilde{D}(p,1)$ is conformal to the band $[0,2\pi]\times[0,\infty)$. Let $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)$ be defined as in (3). Under the conformal identification ${\mathbf{D}_{S}\cap\tilde{D}(p,1)}\approx[0,2\pi]\times[0,\infty)$, there is some deck-transformation $J_{p}$ of $\tilde{S}$ such that $J_{p}((0,h))=(2\pi,h)$ for any $h\geq 0$. Since by conformal invariance, $\operatorname{E}(u|_{[0,2\pi]\times[0,\infty)})=\operatorname{E}(u|_{\mathbf{D}_{S}\cap\tilde{D}(p,1)})<\infty$, by Fubini’s theorem, there are some $h_{j}\to\infty$ such that
$\lim_{j\to\infty}\int_{[0,2\pi]\times\{h_{j}\}}|du|^{2}=0.$
By Cauchy-Schwarz, this implies that
$\lim_{j\to\infty}\operatorname{dist}_{\mathbb{S}_{V}}\big{(}u((0,h_{j})),u((2\pi,h_{j}))\big{)}=0.$
Since $u((2\pi,h_{j}))=u(J_{p}((0,h_{j})))=\rho(J_{p})u((0,h_{j}))$ by equivariance of $u$, it means that $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ for all $p\in\operatorname{Punc}_{S}$. This in turn implies $\operatorname{E}(u|_{\mathbf{D}_{S}})=\operatorname{E_{\mathrm{ren}}}(u|_{\mathbf{D}_{S}})$ by definition of the renormalized energy, so minimizing over $u$, we conclude $\operatorname{E}(S,\rho)=\operatorname{E_{\mathrm{ren}}}(S,\rho).$
We will need the following known monotonicity property of the renormalized energy, see [MRVS22, Lemma 2.11] for instance.
Lemma 1.5.
For any $u\in\mathscr{H}_{S,\rho}$, and any $0<r_{2}<r_{1}\leq 1$,
$\operatorname{E}(u|_{\mathbf{D}_{S}\setminus\bigcup_{p\in\mathrm{Punc}_{S}}\tilde{D}(p,r_{1})})-\sum_{p\in\mathrm{Punc}_{S}}\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p)^{2}}{4\pi}\log\frac{1}{r_{1}}$
$\leq$ $\operatorname{E}(u|_{\mathbf{D}_{S}\setminus\bigcup_{p\in\mathrm{Punc}_{S}}\tilde{D}(p,r_{2})})-\sum_{p\in\mathrm{Punc}_{S}}\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p)^{2}}{4\pi}\log\frac{1}{r_{2}}$
$\leq$ $\operatorname{E_{\mathrm{ren}}}(u|_{\mathbf{D}_{S}}).$
Proof.
Given $p\in\operatorname{Punc}_{S}$, the domain $\mathbf{D}_{S}$ can be chosen so that the closure of $\mathbf{D}_{S}\cap\tilde{D}(p,r_{1})\setminus\tilde{D}(p,r_{2})$ is conformal to the band $[0,2\pi]\times[\log\frac{1}{r_{1}},\log\frac{1}{r_{2}}]$. By definition of $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)$, equivariance of $u$ and Cauchy-Schwarz,
[…]
$\geq\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p)^{2}}{4\pi}\log\frac{1}{r_{2}}-\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p)^{2}}{4\pi}\log\frac{1}{r_{1}}.$
Summing over $p\in\mathrm{Punc}_{S}$, we get the first inequality in the statement. The second inequality follows by letting $r_{2}\to 0$. ∎
1.2. Harmonic maps
Harmonic maps from a surface are smooth maps which are critical points for the energy functional on any compact domain of the surface [SY97]. Geometrically, the main property of the renormalized energy is that, for finite dimensional representations, it is always realized by a harmonic map. Let $S$, $\mathbf{D}_{S}\subset\tilde{S}$, $\rho:\pi_{1}(S)\to\operatorname{End}(V)$, $\mathscr{H}_{S,\rho}$ be as before.
Definition 1.6.
Any map $\psi\in\mathscr{H}_{S,\rho}$ such that
$\operatorname{E_{\mathrm{ren}}}(\psi|_{\mathbf{D}_{S}})=\operatorname{E_{\mathrm{ren}}}(S,\rho).$
is called a harmonic representative of $(S,\rho)$.
Harmonic representatives are clearly harmonic maps. The notion of harmonic representatives does not depend on the specific choice of conformal parameterization of punctured neighborhoods of punctures used to define the renormalized energy. We have the following existence theorem:
Theorem 1.7.
If $\dim_{\mathbb{R}}V<\infty$, then $(S,\rho)$ admits a harmonic representative.
Proof.
We outline the proof, which follows from standard arguments, see [MRVS22, Proposition 8.1] [DM23, Section 4]. Consider a sequence of maps $u_{j}\in\mathscr{H}_{S,\rho}$ whose renormalized energies converge to the infimum $\operatorname{E_{\mathrm{ren}}}(S,\rho)$. We use Lemma 1.5 to bound uniformly the energy of $u_{j}$ on any compact subset of $\mathbf{D}_{S}$ so that classical compactness arguments apply [SY79, SU82] and a susequential limit map exists for each compact subset of $\mathbf{D}_{S}$. After a diagonal argument and taking a union of these limits, we get the desired map. ∎
We will need a “dimension-free” $\varepsilon$-regularity theorem for harmonic maps into spheres:
Theorem 1.8.
For $r>0$, let $\mathbb{D}_{r}$ be the standard disk in $\mathbb{R}^{2}$ of radius $r$. There is a universal constant $\varepsilon_{0}>0$, and for each integer $m\geq 0$ there is a constant $A_{m}>0$ such that the following holds. Let $V$ be a finite dimensional Hilbert space, with unit sphere $S_{V}$. Let $\phi:\mathbb{D}_{r}\to S_{V}$ be a harmonic map with energy at most $\varepsilon_{0}$. Then
$\|\phi\|_{C^{m}(\mathbb{D}_{r/2})}\leq\frac{A_{m}}{r^{m}}.$
Proof.
This result is classical [SU81], except maybe for the fact that the constants $A_{m}$ do not depend on the dimension of $V$. By rescaling and conformal invariance of the energy, it suffices to show the theorem for $r=1$. A bound on $\|\phi\|_{C^{1}(\mathbb{D}_{r^{\prime}})}$ for $r^{\prime}\in(0,1)$, depending only on $r^{\prime}$, can be deduced from the proof in [CM11, pp 149–151], where only a $C^{2}$ bound for the second fundamental form of the sphere $S_{V}$ as a submanifold of $V$ is used, and not its dimension. See [KS20, Lemma 3.12], which made the same observation.
Since we were not able to find a reference, we sketch below how the uniform bounds for the higher derivatives of $\phi$ follow from a bound on $\|\phi\|_{C^{1}(\mathbb{D}_{r^{\prime}})}$ for $r^{\prime}\in(0,1)$. If $n=\dim_{\mathbb{R}}V$, the harmonic map equation is given by
$\Delta\phi^{i}=-|d\phi|^{2}\phi^{i}\quad\text{for all $i=1,...,n$},$
where $\phi^{i}$ are the coordinate functions of $\phi$. Thus, up to a factor depending on $m$, $|d^{m}\Delta\phi^{i}|$ can be bounded by a sum of terms of the form $|d^{m_{0}}\phi^{i}||d^{m_{1}}\phi||d^{m_{2}}\phi|$ where $m_{0},m_{1},m_{2}$ are integers strictly smaller than $m+2$ such that $m_{0}+m_{1}+m_{2}=m+2$. Let $\mathcal{T}(m)$ be the set of such triples of integers $(m_{0},m_{1},m_{2})$. In the remaining of the proof, $C$ denotes a constant (which can change from line to line) depending on certain parameters, but independent of $n$ and $\phi$. By the usual $L^{p}$ estimate [GT77, Theorem 9.11] with $p=2$, for any $m\geq 0$ and any $\delta>0$,
$\int_{\mathbb{D}_{r^{\prime}-\delta}}|d^{m+2}\phi|^{2}$ $=\int_{\mathbb{D}_{r^{\prime}-\delta}}\sum_{i=1}^{n}|d^{m+2}\phi^{i}|^{2}\leq C\sum_{i=1}^{n}\big{(}\int_{\mathbb{D}_{r^{\prime}}}|d^{m}\Delta\phi^{i}|^{2}+\int_{\mathbb{D}_{r^{\prime}}}|d^{m}\phi^{i}|^{2}\big{)}$
$\leq C\sum_{i=1}^{n}\big{(}\int_{\mathbb{D}_{r^{\prime}}}\sum_{(m_{0},m_{1},m_{2})\in\mathcal{T}(m)}(|d^{m_{0}}\phi^{i}||d^{m_{1}}\phi||d^{m_{2}}\phi|)^{2}+\int_{\mathbb{D}_{r^{\prime}}}|d^{m}\phi^{i}|^{2}\big{)}$
$\leq C\sum_{(m_{0},m_{1},m_{2})\in\mathcal{T}(m)}\int_{\mathbb{D}_{r^{\prime}}}|d^{m_{0}}\phi|^{2}|d^{m_{1}}\phi|^{2}|d^{m_{2}}\phi|^{2}+C\int_{\mathbb{D}_{r^{\prime}}}|d^{m}\phi|^{2}.$
We already know that $|\phi|^{2}$ and $|d\phi|^{2}$ are uniformly bounded in $\mathbb{D}_{r^{\prime}}$. Suppose that for an integer $m\geq 2$, $|d^{t}\phi|^{2}\leq C$ on $\mathbb{D}_{r^{\prime}}$ for all $t\leq m-1$, and $\max\{\int_{\mathbb{D}_{r^{\prime}}}|d^{m}\phi|^{2},\int_{\mathbb{D}_{r^{\prime}}}|d^{m+1}\phi|^{2}\}\leq C$, then the inequalities above imply that $\int_{\mathbb{D}_{r^{\prime}-\delta}}|d^{m+2}\phi|^{2}\leq C$ (since $m+(m+1)>m+2$, each factor in $|d^{m_{0}}\phi||d^{m_{1}}\phi||d^{m_{2}}\phi|$ is bounded pointwise by $C$ except maybe one). By the Sobolev inequalities, $|d^{m}\phi|^{2}\leq C$ on $\mathbb{D}_{r^{\prime}-\delta}$. By induction, we deduce that for all $m\geq 0$, $\|\phi\|_{C^{m}(\mathbb{D}_{r^{\prime\prime}})}\leq C$, where $r^{\prime\prime}\in(0,1)$ (the constant $C$ depends on $m,r^{\prime\prime}$ but not on $n,\phi$). This concludes the proof. ∎
1.3. Spherical Plateau problem
The spherical Plateau problem refers to a collection of variational problems depending on an $n$-manifold and an orthogonal representation, and whose solutions (if they exist) are $n$-dimensional minimal surfaces in spheres, invariant under a group action [Son23b, Section 3] [Son24]. In this subsection, we define an invariant relevant to the spherical Plateau problem: the spherical area. This invariant is an important motivation for studying the energy. It generalizes an invariant of Besson-Courtois-Gallot [BCG91, Subsections 3.I and 3.II] to arbitrary unitary representations.
Let $S$ be a punctured Riemann surface of negative Euler characteristic, and let $\Sigma$ be the underlying topological surface. Let $\mathcal{T}_{\Sigma}$ be its Teichmüller space. The Riemann surface $S$ determines a class in Teichmüller space called $[S]$, and vice versa any Teichmüller class is realized by a punctured Riemann surface $S^{\prime}$ with underlying topological surface $\Sigma$. Consider a unitary representation.
$\rho:\pi_{1}(\Sigma)\to\operatorname{End}(H).$
Define $\mathbf{D}_{\Sigma}\subset\tilde{\Sigma}$ and $\mathscr{H}_{\Sigma,\rho}$ in a similar way as for $\mathbf{D}_{S}\subset\tilde{S}$ and $\mathscr{H}_{S,\rho}$.
Definition 1.9.
The spherical area of $(\Sigma,\rho)$ is defined as
$\operatorname{Area}(\Sigma,\rho):=\inf\{\operatorname{E}(S^{\prime},\rho);\quad[S^{\prime}]\in\mathscr{T}_{\Sigma}\}.$
Like the energy, it could be infinite in general. We remark that if $\Sigma$ is a closed surface, then due to classical arguments (see for instance the proof of [SY79, Theorem 3.1]), the spherical area of $(\Sigma,\rho)$ satisfies
$\operatorname{Area}(\Sigma,\rho)=\inf\{\operatorname{Area}(\mathbf{D}_{\Sigma},\phi^{*}g_{V});\quad\phi\in\mathscr{H}_{\Sigma,\rho}\}<\infty.$
Branched minimal surfaces are by definition images of harmonic maps which are weakly conformal. A well-known general strategy to construct minimal surfaces is to divide an area minimization problem into, first, an energy minimization problem inside a conformal class (leading to the construction of a harmonic map), and then, an energy minimization problem on the Teichmüller space of the surface (leading to the desired minimal surface). In our setting, the second step is subtle and will not be treated. We only state the following known fact:
Theorem 1.10.
Let $S$ be a punctured Riemann surface, $\Sigma$ the underlying topological surface and $\rho:\pi_{1}(S)\to\operatorname{End}(H))$ a unitary representation. Suppose that
$\operatorname{E}(S,\rho)=\operatorname{Area}(\Sigma,\rho)<\infty.$
Then, for any map $\psi\in\mathscr{H}_{S,\rho}$ such that $E(\psi|_{\mathbf{D}_{S}})=\operatorname{E}(S,\rho)$, $\psi$ is harmonic and weakly conformal. In other words, $\psi(\tilde{S})$ is a branched minimal surface.
Proof.
This is a corollary of a standard result, which states that a map which is a critical point of the energy with respect to both conformal changes and variations of the conformal class is a branched minimal immersion [SU81, Theorem 1.8].
1.4. Energy and area for the regular representation
Let us describe what happens for the “standard” spherical Plateau problem, where the unitary representation is given by the (left) regular representation. Let $S$ be a punctured Riemann surface of negative Euler characteristic $\chi(S)<0$, let $\Sigma$ be the underlying topological surface.
The regular representation $\lambda_{\pi_{1}(S)}:{\pi_{1}(S)}\to\operatorname{End}(\ell^{2}({\pi_{1}(S)},\mathbb{C}))$ of ${\pi_{1}(S)}$ is the following canonical representation: for all $\gamma,x\in{\pi_{1}(S)}$ and $f\in\ell^{2}({\pi_{1}(S)},\mathbb{C})$,
$(\lambda_{\pi_{1}(S)}(\gamma).f)(x):=f(\gamma^{-1}x).$
It induces a proper free isometric action on the unit sphere $\mathbb{S}_{\ell^{2}({\pi_{1}(S)},\mathbb{C})}$ of the Hilbert space $\ell^{2}({\pi_{1}(S)},\mathbb{C})$. Let $g_{\ell^{2}({\pi_{1}(S)},\mathbb{C})}$ be the standard round metric on $\mathbb{S}_{\ell^{2}({\pi_{1}(S)},\mathbb{C})}$. When $\rho=\lambda_{\pi_{1}(S)}$, the spherical area of $(\Sigma,\rho)$ is exactly the “spherical volume” of $\Sigma$ considered in [BCG91] [Son23b] [Son24]. When $\rho=\lambda_{\pi_{1}(S)}$, the spherical area is also well-defined and easily shown to be equal to 00 for the 2-sphere and the 2-torus. Slightly generalizing Besson-Courtois-Gallot’s computation of the standard spherical area of closed surfaces [BCG91, Remarque 3.1.4 (i)], we have the following:
Theorem 1.11.
Let $\mathbf{g}$ and $\mathbf{n}$ be respectively the genus and the number of punctures of $S$. Then
$\operatorname{E}(S,\lambda_{\pi_{1}(S)})=\operatorname{Area}(\Sigma,\lambda_{\pi_{1}(S)})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)=\frac{\pi}{4}|\chi(S)|.$
Proof.
Let $g_{\mathrm{hyp}}$ be the unique complete finite area hyperbolic metric on $\Sigma$ compatible with the conformal structure of $S$. Note that the number $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)=\frac{\pi}{4}|\chi(S)|$ in Theorem 1.11 is exactly $\operatorname{Area}(\Sigma,\frac{1}{8}g_{\mathrm{hyp}})$.
Let $\phi\in\mathscr{H}_{S,\lambda_{\pi_{1}(S)}}$ and view it as a $\mathbb{C}$-valued function of two variables $x\in\tilde{S}$ and $\gamma\in{\pi_{1}(S)}$. Adapting the proof of [BCG91, Théorème 3.8], we get:
$2E(\phi|_{\mathbf{D}_{S}})$ $=2E(\phi|_{(\mathbf{D}_{\Sigma},g_{\mathrm{hyp}})})$
$=\int_{\mathbf{D}_{\Sigma}}\sum_{\gamma\in{\pi_{1}(S)}}|d_{1}\phi(x,\gamma)|^{2}dv_{g_{\mathrm{hyp}}}(x)=\int_{\mathbf{D}_{\Sigma}}\sum_{\gamma\in{\pi_{1}(S)}}|d_{1}\phi(\gamma^{-1}x,1)|^{2}dv_{g_{\mathrm{hyp}}}(x)$
[…]
The second line holds by ${\pi_{1}(S)}$-equivariance of $\phi$. The inequality comes from the well-known fact that the bottom of the Laplace spectrum for $W^{1,2}$ functions on the hyperbolic plane $(\tilde{\Sigma},g_{\mathrm{hyp}})$ is $\frac{1}{4}$, and the last equality comes from the fact that $\sum_{\gamma\in{\pi_{1}(S)}}|\phi(x,\gamma)|^{2}=1$ for all $x$. Note that we implicitly used the fact that $(\Sigma,g_{\mathrm{hyp}})$ has finite area to ensure that $\phi(.,1)\in W^{1,2}(\tilde{\Sigma})$. Since this computation holds for any $S$ and any $\phi\in\mathscr{H}_{S,\lambda_{\pi_{1}(S)}}$,
(4) $\operatorname{E}(S,\lambda_{\pi_{1}(S)})\geq\operatorname{Area}(\Sigma,\lambda_{\pi_{1}(S)})\geq\operatorname{Area}(S,\frac{1}{8}g_{\mathrm{hyp}}).$
To show the reverse inequalities, for any $c>1$, consider the family of maps
[…]
It is simple to check that $\mathcal{P}_{c}\in\mathscr{H}_{S,\lambda_{\pi_{1}(S)}}$. Furthermore, by [Son23a, Lemma 1.3], for any $x\in\tilde{S}$, and any orthonormal basis $\{e_{1},e_{2}\}$ of $T_{x}\tilde{S}$,
$|d_{x}\mathcal{P}_{c}(e_{1})|^{2}+|d_{x}\mathcal{P}_{c}(e_{2})|^{2}\leq\frac{c^{2}}{4}.$
Thus, by letting $c$ go to $1$, we get
(6) $\operatorname{E}(S,\lambda_{\pi_{1}(S)})\leq\operatorname{Area}(\Sigma,\frac{1}{8}g_{\mathrm{hyp}}).$
Together with (4), this finishes the proof.
The notion of direct sum for Hilbert spaces and unitary representations is defined in [BdLHV08, Definition A.1.6]. Let $\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}$ be the infinite direct sum of $\lambda_{{\pi_{1}(S)}}$. The corresponding Hilbert space is $\bigoplus^{\infty}\ell^{2}({\pi_{1}(S)},\mathbb{C})$. We will later need the following:
Lemma 1.12.
We have
$\operatorname{E}(S,\lambda_{\pi_{1}(S)})=\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})=\operatorname{E}(S,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}})=\operatorname{Area}(\Sigma,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}).$
Proof.
Clearly, since $\lambda_{\pi_{1}(S)}$ is a subrepresentation of $\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}$, we have
$\operatorname{E}(S,\lambda_{\pi_{1}(S)})\geq\operatorname{E}(S,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}})\quad\text{and}\quad\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})\geq\operatorname{Area}(\Sigma,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}).$
Next, we can rewrite $\bigoplus^{\infty}\ell^{2}({\pi_{1}(S)},\mathbb{C})$ as $\ell^{2}({\pi_{1}(S)},H)$ where $H$ is the separable infinite dimensional Hilbert space. Define the following map between the unit spheres
$\mathcal{A}:\mathbb{S}_{\ell^{2}({\pi_{1}(S)},H)}\to\mathbb{S}_{\ell^{2}({\pi_{1}(S)},\mathbb{C})}$
such that for any $f\in\mathbb{S}_{\ell^{2}({\pi_{1}(S)},H)}$ and $\gamma\in{\pi_{1}(S)}$, $\mathcal{A}(f)(\gamma):=|f(\gamma)|_{H}$ where |.$|.|_{H}$ denotes the norm in $H$. Clearly, $\mathcal{A}$ is distance non-increasing and equivariant. If $\phi\in\mathscr{H}_{\Sigma,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}}$, then $\mathcal{A}\circ\phi$ is still equivariant. For any $\epsilon>0$, after smoothing $\mathcal{A}\circ\phi$, we get a map $\phi^{\prime}\in\mathscr{H}_{S,\lambda_{{\pi_{1}(S)}}}$ such that $E(\phi^{\prime}|_{\mathbf{D}_{S}})\leq E(\phi|_{\mathbf{D}_{S}})+\epsilon$ and so
$\operatorname{E}(S,\lambda_{\pi_{1}(S)})\leq\operatorname{E}(S,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}).$
Taking the infimum over conformal classes $[S]\in\mathcal{T}_{\Sigma}$, this implies
$\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})\leq\operatorname{Area}(\Sigma,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}}).$
Next, we consider a special unitary representation and a special embedding $\mathscr{P}$ of the hyperbolic plane into a Hilbert sphere, closely related to the maps $\mathcal{P}_{c}$ defined earlier in (5), and which played a crucial role in Besson-Courtois-Gallot’s paper on the entropy inequality [BCG95] (see [Son23b, Subsection 4.2]). First, if $g_{\mathrm{hyp}}$ is as usual the unique conformal hyperbolic metric on $S$, we view $\pi_{1}(S)$ as a subgroup of the oriented isometry group $\mathrm{PSL_{2}(\mathbb{R})}$ of the hyperbolic plane $(\tilde{S},g_{\mathrm{hyp}})$. Fix a basepoint $\mathbf{o}\in\tilde{S}$, let $\partial\tilde{S}$ be the boundary at infinity of $\tilde{S}$ with the standard uniform probability measure determined by $\mathbf{o}$, and let $\mathbb{S}_{2}(\partial\tilde{S})$ be the unit sphere in the complex $L^{2}$-space $L^{2}(\partial\tilde{S})$ endowed with the standard Riemannian metric $g_{L^{2}(\partial\tilde{S})}$. For any $\theta\in\partial\tilde{S}$, the corresponding Busemann function is defined for any $x\in\tilde{S}$ as
$B_{\theta}(x):=\lim_{t\to\infty}(\operatorname{dist}_{g_{\mathrm{hyp}}}(y,c(t))-t)$
where $c:[0,\infty)$ is the half-geodesic starting at $\mathbf{o}$, and converging to $\theta$. The group ${\pi_{1}(S)}$ acts naturally on $\tilde{S}$ and $\partial\tilde{S}$ as a subgroup of $\mathrm{PSL_{2}(\mathbb{R})}$. This induces a unitary representation, called “boundary representation”:
$\underline{\rho}_{B}:{\pi_{1}(S)}\to\operatorname{End}(L^{2}(\partial\tilde{S}))$
which is defined, for all $\gamma\in{\pi_{1}(S)}$, $f\in\mathbb{S}_{2}(\partial\tilde{S})$, by
(7) $\underline{\rho}_{B}(\gamma).f(\theta)=f(\gamma^{-1}(\theta))e^{-\frac{1}{2}B_{\theta}(\gamma(\mathbf{o}))},$
see [BCG95, Lemme 2.2]. Next, consider the following embedding
(8) $\begin{split}\mathscr{P}:(\tilde{S},g_{\mathrm{hyp}})&\to(\mathbb{S}_{2}(\partial\tilde{S}),g_{L^{2}(\partial\tilde{S})})\\ \mathscr{P}(x)&:=\{\theta\mapsto e^{-\frac{1}{2}B_{\theta}(x)}\}.\end{split}$
As explained in [BCG95, Section 2], this map is ${\pi_{1}(S)}$-equivariant, namely it belongs to $\mathscr{H}_{S,\underline{\rho}_{B}}$. The notions of weak equivalence and weak containment for representations are recalled in Definition 6.1 in the Appendix. We need the following facts about $\underline{\rho}_{B}$ and $\mathscr{P}$:
Lemma 1.13.
  1. (1)
    The boundary representation $\underline{\rho}_{B}$ is irreducible, and weakly equivalent to the regular representation $\lambda_{\pi_{1}(S)}$.
  2. (2)
    The map $\mathscr{P}$ is an isometric embedding of the rescaled hyperbolic plane $(\tilde{S},\frac{1}{8}g_{\mathrm{hyp}})$ into $\mathbb{S}_{2}(\partial\tilde{S})$. In particular,
    $\operatorname{E}(\mathscr{P}|_{\mathbf{D}_{S}})=\operatorname{Area}(\Sigma,\lambda_{\pi_{1}(S)})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2).$
Proof.
(1) The group ${\pi_{1}(S)}$ is a lattice of the simple Lie group $\mathrm{PSL}_{2}(\mathbb{R})$ of isometries of the hyperbolic plane. The representation $\underline{\rho}_{B}$ is the restriction of a unitary representation $\hat{\rho}_{B}:\mathrm{PSL}_{2}(\mathbb{R})\to\operatorname{End}(L^{2}(\partial\tilde{S}))$ defined by the generalization of formula (7) to the whole Lie group $\mathrm{PSL}_{2}(\mathbb{R})$. Actually, $\hat{\rho}_{B}$ can be identified with the irreducible representation from the principal series called $\rho_{0}$ in the notation of [Lub94, Theorem 5.2.3], where the author reviews the classification of irreducible representations of $\mathrm{PSL}_{2}(\mathbb{R})$ containing $SO(2)/\{\pm 1\}$-invariant vectors. Indeed, the representation $\hat{\rho}_{B}$ is cyclic ([BdLHV08, Definition C.4.8]): for instance it can be easily checked that for any nonzero $f\in L^{2}(\partial\tilde{S})$, there is some $g\in\mathrm{PSL}_{2}(\mathbb{R})$ such that $\langle\hat{\rho}_{B}(g)1,f\rangle_{L^{2}(\partial\tilde{S})}\neq 0$. Moreover, the constant function $1$ in $L^{2}(\partial\tilde{S})$ is fixed by $SO(2)/\{\pm 1\}$, and
$\varphi:g\mapsto\langle\hat{\rho}_{B}(g)1,1\rangle_{L^{2}(\partial\tilde{S})}$
is a spherical function in the sense of [Lub94, Definition 5.1.5]. As a function on the hyperbolic plane, by a direct computation using (7), it satisfies the equation $\Delta\varphi=-\frac{1}{4}\varphi$. Since this spherical function uniquely determines the cyclic representation $\hat{\rho}_{B}$ by the GNS construction [BdLHV08, Theorem C.4.10], we can identify $\hat{\rho}_{B}$ with $\rho_{0}$ from the principal series by [Lub94, Definition 5.1.5], as claimed. In particular, $\hat{\rho}_{B}$ is not a discrete series, since those do not contain $SO(2)/\{\pm 1\}$-invariant vectors. The irreducibility of $\underline{\rho}_{B}$ is then a consequence of [SC91, Proposition 2.5 (ii)].
By $C^{*}$-simplicity of surface groups such as $\pi_{1}(S)$ [dLH07, Definitions and examples on pages 2 and 3], in order to show that $\underline{\rho}_{B}$ is weakly equivalent to $\lambda_{\pi_{1}(S)}$, it suffices to show that the first is weakly contained in the second. The fact that $\underline{\rho}_{B}$ is weakly contained in $\lambda_{\pi_{1}(S)}$ follows from the classical property that $\hat{\rho}_{B}$ is weakly contained in $\lambda_{\mathrm{PSL}_{2}(\mathbb{R})}$, or can be checked directly by “approximating by hand” functions in $L^{2}(\partial\tilde{S})$ by functions in $\ell^{2}({\pi_{1}(S)},\mathbb{C})$.
(2) Checking that $\mathscr{P}$ is an isometry is a computation done in [BCG95, 2.6 Retour aux exemples. a)] (translating their notations to our case, $n=2$, $h_{0}=1$, $g_{\sqrt{p_{0}}}=\mathscr{P}^{*}g_{L^{2}(\partial\tilde{S})}$).
Remark 1.14.
As we will see later, the relevance of the special embedding $\mathscr{P}$ is that it models the limit of equivariant harmonic maps into finite-dimensional spheres, which are equivariant with respect to strongly converging representations. What are the other possible limits of equivariant harmonic maps into finite-dimensional spheres?
2. Intrinsic uniqueness of equivariantly minimizing surfaces
The goal of this subsection is to prove a general intrinsic uniqueness result for equivariant minimal surfaces in spheres, which seems to be the first rigidity result in this spirit. This will play a key role: we will use it to identify the unique limit of the pullback metric by the harmonic maps of the main theorem.
Let $S$ be a punctured Riemann surface, with universal cover $\tilde{S}$, and fundamental domain $\mathbf{D}_{S}$. Let $g$ be any Riemannian metric on $S$ compatible with its conformal structure. Let $\Sigma$ be the topological surface underlying $S$. Consider a unitary representation $\rho:{\pi_{1}(S)}\to\operatorname{End}(H^{\prime})$ and its infinite direct sum (see [BdLHV08, Definition A.1.6]):
$\bigoplus^{\infty}\rho:{\pi_{1}(S)}\to\operatorname{End}(\bigoplus^{\infty}H^{\prime}).$
The main result of this subsection is that in the conformal class of $S$, there is at most one Riemannian metric $g_{1}$ for which there exists an equivariant, branched, minimal, isometric immersion from $(\tilde{S},g_{1})$ to the unit sphere $\mathbb{S}_{\bigoplus^{\infty}H^{\prime}}$, which achieves $\operatorname{Area}(\Sigma,\bigoplus^{\infty}\rho)$. We take the infinite direct sum of $\rho$, because the “self-similarity” of $\bigoplus^{\infty}\rho$ turns out to be an extremely helpful feature of the infinite dimensional setting. [^5] The result probably fails for $\rho$ in place of $\bigoplus^{\infty}\rho$.
Theorem 2.1.
Consider two functions $f_{1},f_{2}:\tilde{S}\to[0,\infty)$ which are allowed to vanish at a discrete set of points. Suppose that for each $i=1,2$, there is a branched, minimal, isometric immersion
$\varphi_{i}:(\tilde{S},f_{i}^{2}g)\to\mathbb{S}_{\bigoplus^{\infty}H^{\prime}}$
which is ${\pi_{1}(S)}$-equivariant with respect to $\bigoplus^{\infty}\rho$ and such that
$\operatorname{E}(\varphi_{i}|_{\mathbf{D}_{S}})=\operatorname{Area}(\Sigma,\bigoplus^{\infty}\rho).$
Then $f_{1}=f_{2}$.
Remark 2.2.
The meaning of Theorem 2.1 is that surfaces in spheres which are equivariant with respect to a “self-similar” representation and equivariantly area-minimizing are intrinsically unique in each conformal class (if they exist). Are they also extrinsically unique?
Theorem 2.1 will directly follow from the next lemma. Let $H$, $K$ be two Hilbert spaces, whose direct sum is denoted by $H\oplus K$ and is endowed with the direct sum of the scalar products of $H$, $K$. Denote their respective unit spheres by $\mathbb{S}_{H}$, $\mathbb{S}_{K}$, $\mathbb{S}_{H\oplus K}$. Consider two conformal factors $f_{1},f_{2}:\tilde{S}\to[0,\infty)$ which are allowed to vanish at a discrete set of points, and the two conformal metrics $f_{1}^{2}g$, $f_{2}^{2}g.$ Suppose that there are two minimal isometric branched immersions
$\varphi_{1}:(\tilde{S},f_{1}^{2}g)\to\mathbb{S}_{H},\quad\varphi_{2}:(\tilde{S},f_{2}^{2}g)\to\mathbb{S}_{K}.$
By the classical result of Takahashi  (adapted to minimal surfaces in Hilbert spheres) [Tak66], and by conformal invariance of the Laplacian in dimension $2$, we have around points where $f_{1}\neq 0,f_{2}\neq 0$:
(9) $\Delta_{g}\varphi_{1}=-2f_{1}^{2}\varphi_{1},\quad\Delta_{g}\varphi_{2}=-2f_{2}^{2}\varphi_{2}.$
Define the “average” branched immersion of $\varphi_{1},\varphi_{2}$ to be
$\varphi_{3}:\tilde{S}\to\mathbb{S}_{H\oplus K},$
$\varphi_{3}(x):=\frac{1}{\sqrt{2}}(\varphi_{1}(x)\oplus\varphi_{2}(x)).$
Note that $\varphi_{3}$ is automatically a conformal map, and that
(10) $\operatorname{E}(\varphi_{3}|_{D_{S}})=\frac{1}{2}(\operatorname{E}(\varphi_{1}|_{D_{S}})+\operatorname{E}(\varphi_{2}|_{D_{S}})).$
Lemma 2.3.
If the average branched immersion $\varphi_{3}$ is harmonic, then $f_{1}=f_{2}$.
Proof.
If $\varphi_{3}$ is harmonic then
$\Delta_{g}\varphi_{3}=a_{3}\varphi_{3}$
for some function $a_{3}:\tilde{S}\to\mathbb{R}$. On the other hand, by definition of $\varphi_{3}$, and by (9):
$\Delta_{g}\varphi_{3}=\frac{1}{\sqrt{2}}(\Delta_{g}\varphi_{1}\oplus\Delta_{g}\varphi_{2})=-\sqrt{2}(f_{1}^{2}\varphi_{1}\oplus f_{2}^{2}\varphi_{2}).$
In order for these two equations to be compatible, necessarily ${f_{1}}={f_{2}}$.
Proof of Theorem 2.1.
Set $H=K=\bigoplus^{\infty}H^{\prime}$. A crucial remark is that $\bigoplus^{\infty}\rho$ is self-similar, in the sense that the representation
(11) $(\bigoplus^{\infty}\rho\oplus\bigoplus^{\infty}\rho,H\oplus K)$
is equivalent to $(\bigoplus^{\infty}\rho,H)$. Under our assumptions, the average $\varphi_{3}$ of $\varphi_{1},\varphi_{2}$ is a branched conformal immersion
$\varphi_{3}:\tilde{S}\to\mathbb{S}_{H\oplus K}\approx\mathbb{S}_{H}$
which is ${\pi_{1}(S)}$-equivariant with respect to $\bigoplus^{\infty}\rho\oplus\bigoplus^{\infty}\rho$, and by (10), its energy necessarily satisfies
$\operatorname{E}(\varphi_{3}|_{D_{S}})=\operatorname{Area}(\Sigma,\bigoplus^{\infty}\rho).$
By (11) and definition of the spherical area and energy, this implies that $\varphi_{3}$ is also harmonic. Applying Lemma 2.3, we conclude the proof. ∎
We will need the following consequence of Theorem 2.1 for the regular representation, which has a higher-dimensional analogue [Son23b, Corollary 4.3] (proved using completely different methods based on the barycenter map [BCG95]). Let $\mathbf{g}$ and $\mathbf{n}$ be the genus and number of punctures of $S$, and let $g_{\mathrm{hyp}}$ denote the unique complete, finite area, conformal hyperbolic metric on the Riemannian surface $S$ or its lift to $\tilde{S}$. The notions of weak equivalence and weak containment for representations, denoted with the symbols $\sim$ and $\prec$ respectively, are recalled in Definition 6.1 in the Appendix.
Corollary 2.4.
Let $\rho_{1}:{\pi_{1}(S)}\to\operatorname{End}(H_{1})$ be a unitary representation with $\rho_{1}\sim\lambda_{{\pi_{1}(S)}}$ and let $(\mathbb{S}_{H_{1}},g_{H_{1}})$ be the unit sphere in $H_{1}$ with its standard Riemannian metric. Consider a smooth ${\pi_{1}(S)}$-equivariant map
$\varphi_{1}:\tilde{S}\to(\mathbb{S}_{H_{1}},g_{H_{1}}).$
Then
$\operatorname{E}(\varphi_{1}|_{\mathbf{D}_{S}})\geq\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)=\frac{\pi}{4}|\chi(S)|$
and equality holds if and only if
$\varphi_{1}^{*}g_{H_{1}}=\frac{1}{8}g_{\mathrm{hyp}}.$
Proof.
The value of $\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})$ is computed in Theorem 1.11. Consider the unitary representation
$\rho:=\rho_{1}\oplus\underline{\rho}_{B}:{\pi_{1}(S)}\to\operatorname{End}(H_{1}\oplus L^{2}(\partial\tilde{S}))$
where $\underline{\rho}_{B}$ is the boundary representation defined in (7), and
$H:=H_{1}\oplus L^{2}(\partial\tilde{S}).$
Note that since $\rho_{1}\prec\lambda_{\pi_{1}(S)}$ and $\underline{\rho}_{B}\prec\lambda_{\pi_{1}(S)}$,
$\bigoplus^{\infty}\rho\prec\lambda_{{\pi_{1}(S)}}.$
We have by definition
(12) $\operatorname{E}(S,\bigoplus^{\infty}\rho)\leq\operatorname{E}(\varphi_{1}|_{\mathbf{D}_{S}}).$
Thanks to Corollary 3.4 (1) which we will prove in Section 3, we obtain
(13) $\operatorname{E}(S,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}})\leq\operatorname{E}(S,\bigoplus^{\infty}\rho).$
But from Lemma 1.12 and Theorem 1.11,
$\operatorname{E}(S,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}})$ $=\operatorname{Area}(\Sigma,\bigoplus^{\infty}\lambda_{{\pi_{1}(S)}})=\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})=\operatorname{E}(S,\lambda_{{\pi_{1}(S)}})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2).$
Hence, by (12), (13) and the line above, we already obtain the inequality in the statement of the corollary modulo results in Section 3.
Suppose that equality holds: $\operatorname{E}(\varphi_{1}|_{\mathbf{D}_{S}})=\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})$. Together with (12) and (13), we deduce
$\operatorname{E}(\varphi_{1}|_{\mathbf{D}_{S}})=\operatorname{Area}(\Sigma,\lambda_{{\pi_{1}(S)}})=\operatorname{Area}(\Sigma,\bigoplus^{\infty}\rho).$
So by Theorem 1.10, $\varphi_{1}$ is both harmonic and weakly conformal, namely a branched minimal immersion. Consider a fixed Riemannian metric $g$ on $S$ compatible with its conformal structure, and the conformal factor $f_{1}:\tilde{S}\to[0,\infty)$, which vanishes only at a discrete set of points, such that
$\varphi_{1}^{*}g_{H_{1}}=f_{1}^{2}g$
where $g$ is the lift of the metric on $S$. Consider the branched minimal immersion $\varphi_{1}$ and the special minimal embedding $\mathscr{P}$ defined in (8), equivariant with respect to $\underline{\rho}_{B}$. By Lemma 1.13 (2), both have the same energy on the fundamental domain, and $\mathscr{P}$ is an isometric embedding of the rescaled hyperbolic plane $(\tilde{S},\frac{1}{8}g_{\mathrm{hyp}})$ where $g_{\mathrm{hyp}}$ is the lift of the unique hyperbolic metric on $S$ conformal to $g$. Applying the intrinsic uniqueness property in Theorem 2.1, we deduce that $f_{1}^{2}g=\frac{1}{8}g_{\mathrm{hyp}}$ and the corollary is proved. ∎
3. Bounds on the renormalized energy
This section extracts the geometric consequences of notions from representation theory like weak containment. Some of the statements are quite technical but the main proof ideas are simple.
3.1. Weak containment and approximations of equivariant maps
In this subsection, we discuss how the notion of weak containment of representations relates to geometric approximations of maps from compact regions of the universal cover of a surface.
Let $S$ be a punctured Riemann surface. Let $\tilde{S}$ be its universal cover, $\mathbf{D}_{S}$ a Borel fundamental domain in $\tilde{S}$ which can be chosen to be a domain with piecewise smooth boundary. Let $\mathrm{Punc}_{S}=\{p_{1},...,p_{n}\}$ be the possibly empty set of punctures of $S$.
If $\mathbf{n}>0$, recall from Subsection 1.1 that there are disjoint punctured disks
$D(p_{1},1),...,D(p_{\mathbf{n}},1)\subset S$
such that for each $q\in\{1,...,\mathbf{n}\}$, there is a conformal diffeomorphism
$P_{p_{q}}:D(p_{q},1)\to\mathbb{D}^{*},$
and that $D(p_{q},r)$ is defined as the set of points in $D(p_{q},1)$ sent by $P_{p_{q}}$ to points $r$-close to the origin in $\mathbb{D}^{*}$. For any $\delta\in(0,1]$, set
(14) $A_{\delta}:=S\setminus\bigcup_{q=1}^{\mathbf{n}}D(p_{q},\delta).$
If $\mathbf{n}=0$, define by convention $A_{\delta}=S$. Let $\tilde{A}_{\delta}$ be the lift of $A_{\delta}$ to $\tilde{S}$.
If $H$ is a Hilbert space, we will always denote by $g_{H}$ its Hilbert metric and $\mathbb{S}_{H}$ its unit sphere. Given a unitary representation $\nu:{\pi_{1}(S)}\to\operatorname{End}(K)$, as before we denote by $\bigoplus^{\infty}\nu:{\pi_{1}(S)}\to\operatorname{End}(\bigoplus^{\infty}K)$ the infinite direct sum ([BdLHV08, Definition A.1.6]). For the definitions of weak containment and weak equivalence of representations, and the corresponding symbols $\prec$ and $\sim$, see Subsection 6.1 in the Appendix.
Proposition 3.1.
Let $\delta\in(0,1]$. Consider two unitary representations $\rho:{\pi_{1}(S)}\to\operatorname{End}(H)$ and $\nu:{\pi_{1}(S)}\to\operatorname{End}(K)$. Suppose that $\rho\prec\nu$. Let
$\varphi:\tilde{A}_{\delta}\to\mathbb{S}_{H}$
be a smooth ${\pi_{1}(S)}$-equivariant map with respect to $\rho$. Then for any $\epsilon^{\prime}>0$,
  1. (1)
    there is a smooth ${\pi_{1}(S)}$-equivariant map with respect to $\bigoplus^{\infty}\nu$,
    $\varphi^{\prime}:\tilde{A}_{\delta}\to\mathbb{S}_{\bigoplus^{\infty}K}$
    such that
    $\|(\varphi^{\prime})^{*}g_{\bigoplus^{\infty}K}-\varphi^{*}g_{H}\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}<\epsilon^{\prime}$
    and
    $|\operatorname{E}(\varphi^{\prime}|_{\mathbf{D}_{S}})-\operatorname{E}(\varphi|_{\mathbf{D}_{S}})|<\epsilon^{\prime},$
  2. (2)
    if moreover $\rho$ is irreducible, then for any direct sum decomposition
    $K=\bigoplus_{m=1}^{\infty}K_{m},\quad\nu=\bigoplus_{m=1}^{\infty}\nu_{m},\quad\nu_{m}:{\pi_{1}(S)}\to\operatorname{End}(K_{m}),$
    the map $\varphi^{\prime}$ can be chosen to take values in $\mathbb{S}_{K_{m_{0}}}$ for some $m_{0}$ depending on $\rho,\epsilon^{\prime}$.
Proof.
It is possible to find a finite or infinite sequence of unit vectors $\{e_{1},e_{2},e_{3},...\}$ such that the family of vectors $\{\rho(g)e_{l}\}_{l\geq 1,g\in{\pi_{1}(S)}}$ generates $H$, and for any $l_{1}\neq l_{2}$, the subspaces generated by $\{\rho(g)e_{l_{1}}\}_{g\in{\pi_{1}(S)}}$ and $\{\rho(g)e_{l_{2}}\}_{g\in{\pi_{1}(S)}}$ are orthogonal.
We fix an ordering on ${\pi_{1}(S)}$. For each $l\geq 1$, let $\{u_{l,j}\}_{j\geq 1}$ be the unique orthonormal Hilbert basis of the Hilbert subspace generated by $\{\rho(g)e_{l}\}_{g\in{\pi_{1}(S)}}$ given by the Gram-Schmidt algorithm applied to $\{\rho(g)e_{l}\}_{g\in{\pi_{1}(S)}}$ using the ordering on ${\pi_{1}(S)}$.
We can choose the fundamental domain $\mathbf{D}_{S}$ so that the compact domain $\overline{\mathbf{D}}_{S}\cap\tilde{A}_{\delta}$ has a piecewise smooth boundary, where $\overline{\mathbf{D}}_{S}$ is the closure of $\mathbf{D}_{S}$. The restriction of $\varphi$ to $\mathbf{D}_{S}\cap\tilde{A}_{\delta}$ can be written under the form:
$\varphi:\mathbf{D}_{S}\cap\tilde{A}_{\delta}\to\mathbb{S}_{H}$
$x\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}\mapsto\sum_{l\geq 1}\sum_{j\geq 1}h_{l,j}(x)u_{l,j}$
for some uniquely determined functions $h_{l,j}:\mathbf{D}_{\Sigma}\cap\tilde{A}_{\delta}\mapsto\mathbb{R}$ which vary smoothly with respect to $x$.
Let $\epsilon>0$. By a standard approximation argument, we can find functions
$h^{\epsilon}_{g}:\mathbf{D}_{S}\cap\tilde{A}_{\delta}\mapsto\mathbb{R}$
such that the new map
$\varphi^{\epsilon}:\mathbf{D}_{S}\cap\tilde{A}_{\delta}\to\mathbb{S}_{H}$
$x\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}\mapsto\sum_{l\geq 1}\sum_{j\geq 1}h^{\epsilon}_{l,j}(x)u_{l,j}$
satisfies two properties:
  1. (1)
    $h^{\epsilon}_{l,j}(x)=0$ for any $x\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}$, except for finitely many $l\geq 1$ and $j\geq 1$,
  2. (2)
    $\|(\varphi^{\epsilon})^{*}g_{H}-\varphi^{*}g_{H}\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}\leq\epsilon.$
In particular, for any $\epsilon^{\prime}>0$, if $\epsilon$ is small enough, then the maps above satisfy:
$\|(\varphi^{\epsilon})^{*}g_{H}-\varphi^{*}g_{H}\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}\leq\epsilon^{\prime}/3,$
$|\operatorname{E}(\varphi^{\epsilon}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})-\operatorname{E}(\varphi|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})|<\epsilon^{\prime}/3.$
Since $\{u_{l,j}\}_{j\geq 1}$ was obtained by applying the Gram-Schmidt procedure to $\{\rho(g)e_{l}\}_{g\in{\pi_{1}(S)}}$, the functions $h^{\epsilon}_{l,j\geq 1}$ determine uniquely a family of functions
$\{k^{\epsilon}_{l,g}:\mathbf{D}_{S}\cap\tilde{A}_{\delta}\mapsto\mathbb{R}\}_{l\geq 1,g\in{\pi_{1}(S)}}$
such that for any $x\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}$,
$\varphi^{\epsilon}(x)=\sum_{l\geq 1}\sum_{g\in{\pi_{1}(S)}}k^{\epsilon}_{l,g}(x)\rho(g)e_{l}$
and for all but finitely many $l\geq 1$ and $g\in{\pi_{1}(S)}$, we have $k^{\epsilon}_{l,g}(x)=0$ for any $x$.
By our assumption that $\rho\prec\nu$, for any $l\geq 0$, for all $\epsilon_{1}$ and every finite subset $Q\subset{\pi_{1}(S)}$, there exist $\eta_{l,1},...,\eta_{l,m}$ in $K$ such that for all $g\in Q$,
$\big{|}\langle\rho(g)e_{l},e_{l}\rangle-\sum_{j=1}^{m}\langle\nu(g)\eta_{l,j},\eta_{l,j}\rangle\big{|}<\epsilon_{1}.$
We can reformulate that by saying the following: for any $l\geq 0$, for all $\epsilon_{1}$ and every finite subset $Q\subset{\pi_{1}(S)}$, there exist $\eta_{l}$ in $\bigoplus^{\infty}K$ such that for all $g\in Q$,
(15) $\big{|}\langle\rho(g)e_{l},e_{l}\rangle-\langle\bigoplus^{\infty}\nu(g)\eta_{l},\eta_{l}\rangle\big{|}<\epsilon_{1}.$
Moreover, if $l_{1}\neq l_{2}$, then the subspace generated by $\eta_{l_{1}}$ is orthogonal to the one generated by $\eta_{l_{2}}$ inside $\bigoplus^{\infty}K$. Inspecting the inequality for $g=e$ the unit element, we can assume that each $\eta_{l}$ has norm one.
We choose the finite subset $Q\subset{\pi_{1}(S)}$ so that it contains all elements of the form $g_{1}^{-1}g_{2}$ where $g_{1},g_{2}\in{\pi_{1}(S)}$ are such that $k^{\epsilon}_{l,g_{1}}(x_{1})\neq 0,k^{\epsilon}_{l,g_{2}}(x_{2})\neq 0$ for some $l\geq 1$ and some $x_{1},x_{2}\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}$. Given $\epsilon_{1}>0$, let $\eta_{l}\in\bigoplus^{\infty}K$ be as above. Define the approximation map
$\mathscr{A}^{\epsilon}:\mathbf{D}_{S}\cap\tilde{A}_{\delta}\to\mathbb{S}_{\bigoplus^{\infty}K}$
$x\in\mathbf{D}_{S}\cap\tilde{A}_{\delta}\mapsto\frac{1}{\|\sum_{l\geq 1}\sum_{g\in{\pi_{1}(S)}}k^{\epsilon}_{l,g}(x)\bigoplus^{\infty}\nu(g)\eta_{l}\|}\sum_{l\geq 1}\sum_{g\in{\pi_{1}(S)}}k^{\epsilon}_{l,g}(x)\bigoplus^{\infty}\nu(g)\eta_{l}.$
By (15) and using that the unitary representations $\rho$ and $\bigoplus^{\infty}\nu$ preserve scalar products, this map is well-defined for $\epsilon_{1}$ small enough since then, the vector on the right side above is nowhere vanishing. By (15) again, for any $\epsilon^{\prime}$, if $\epsilon_{1}$ is chosen small enough, then $\mathscr{A}^{\epsilon}$ and $\varphi^{\epsilon}$ satisfy:
$\|(\mathscr{A}^{\epsilon})^{*}g_{\bigoplus^{\infty}K}-(\varphi^{\epsilon})^{*}g_{H}\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}\leq\epsilon^{\prime}/3,$
$|\operatorname{E}(\mathscr{A}^{\epsilon}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})-\operatorname{E}(\varphi^{\epsilon}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})|<\epsilon^{\prime}/3.$
It remains to construct from $\mathscr{A}^{\epsilon}$ a nearby map that is the restriction to $\mathbf{D}_{S}\cap\tilde{A}_{\delta}$ of a ${\pi_{1}(S)}$-equivariant map from $\tilde{A}_{\delta}$ to $\mathbb{S}_{\bigoplus^{\infty}K}$. Let us call $s_{1},...,s_{j},...$ the finitely many sides of the polygon $\overline{\mathbf{D}}_{S}\cap\tilde{A}_{\delta}$. For any side $s_{j}$ of $\overline{\mathbf{D}}_{S}\cap\tilde{A}_{\delta}$, if $g.s_{j}$ is another side $s_{k}$ of $\overline{\mathbf{D}}_{S}\cap\tilde{A}_{\delta}$ for some $g\in{\pi_{1}(S)}$, then by ${\pi_{1}(S)}$-equivariance of $\varphi$, the two maps $\varphi\circ g:s_{j}\to\mathbb{S}_{H}$ and $\rho(g)\circ\varphi:s_{j}\to\mathbb{S}_{H}$ are equal. It is then clear that if $\epsilon$, $\epsilon_{1}$ are small enough, $\mathscr{A}^{\epsilon}\circ g:s_{j}\to\mathbb{S}_{\bigoplus^{\infty}K}$ and $\bigoplus^{\infty}\nu(g)\circ\mathscr{A}^{\epsilon}:s_{j}\to\mathbb{S}_{\bigoplus^{\infty}K}$ are arbitrarily close in the $C^{\infty}$-topology. One can make them equal by a small perturbation of $\mathscr{A}^{\epsilon}$. Repeating this for every side $s_{j}$, it is then quite elementary to perturb the map $\mathscr{A}^{\epsilon}$ to get a new map
$\mathscr{A}^{\epsilon}_{0}:\overline{\mathbf{D}}_{S}\cap\tilde{A}_{\delta}\to\mathbb{S}_{\bigoplus^{\infty}K}$
such that $\mathscr{A}^{\epsilon}_{0}$ is the restriction of a smooth ${\pi_{1}(S)}$-equivariant map
$\varphi^{\prime}:\tilde{A}_{\delta}\to\mathbb{S}_{\bigoplus^{\infty}K},$
and
$\|(\varphi^{\prime})^{*}g_{\bigoplus^{\infty}K}-(\mathscr{A}^{\epsilon})^{*}g_{\bigoplus^{\infty}K}\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}\leq\epsilon^{\prime}/3,$
$|\operatorname{E}(\varphi^{\prime}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})-\operatorname{E}(\mathscr{A}^{\epsilon}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})|<\epsilon^{\prime}/3.$
Combining the previous inequalities, we get
$\|\varphi^{\prime}-\varphi\|_{C^{3}(\mathbf{D}_{S}\cap\tilde{A}_{\delta})}\leq\epsilon^{\prime},$
$|\operatorname{E}(\varphi^{\prime}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})-\operatorname{E}(\varphi|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})|<\epsilon^{\prime}.$
Property (1) of the proposition is proved.
To show (2), note that if $\rho$ is irreducible, then $H$ is generated by $\{\rho(g)e_{1}\}_{g\in{\pi_{1}(S)}}$ (i.e. there is only one nonzero vector $e_{l}$). Moreover, since $\rho\prec\nu$, if $K=\bigoplus_{m=1}^{\infty}K_{m}$, $\nu=\bigoplus_{m=1}^{\infty}\nu_{m}$, then (15) holds with $\eta_{1}$ belonging to some $K_{m}$ instead of $\bigoplus^{\infty}K$. Indeed, this follows from the proof of [BdLHV08, Proposition F.1.4] (in its proof, replace $\mathcal{F}$ by the set of normalized functions of positive type on ${\pi_{1}(S)}$ of the form $\langle\nu(.)\xi,\xi\rangle$ where $\xi\in\bigcup_{m=1}^{\infty}K_{m}$). Thus, the approximations $\mathscr{A}^{\epsilon}$ and $\mathscr{A}^{\epsilon}_{0}$ can be chosen to take values in $\mathbb{S}_{K_{m}}$ instead of $\mathbb{S}_{\bigoplus^{\infty}K}$, and so $\varphi^{\prime}$ can be chosen to take values in $\mathbb{S}_{K_{m}}$ instead of $\mathbb{S}_{\bigoplus^{\infty}K}$.
3.2. Extending equivariant maps
Below, we explain how to equivariantly extend a map from a compact subset of a universal cover (like those produced in the previous subsection) to the whole universal cover, in such a way that the renormalized energy on a fundamental domain remains controlled. A necessary and sufficient condition for a pair $(S,\rho)$ to have finite (non-renormalized) energy is also given in this subsection.
Let $S$ be a punctured Riemann, with a fundamental domain $\mathbf{D}_{S}\subset\tilde{S}$ as before, and let $\mathrm{Punc}_{S}$ be the set of punctures $p_{1},...,p_{\mathbf{n}}$ of $S$. Let $\rho:{\pi_{1}(S)}\to\operatorname{End}(V)$ be a unitary representation (possibly of infinite dimension), $\mathbb{S}_{V}$ the unit sphere of the Hilbert space $V$. Let $\mathscr{H}_{S,\rho}$ be the space of smooth equivariant maps defined in (1). Given $p\in\mathrm{Punc}_{S}$, the quantity $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)$ was defined in (3).
Preparation and some notations: Let $\delta\in(0,1]$. If $\mathbf{n}>0$, let $A_{\delta}\subset S$ and $\tilde{A}_{\delta}\subset\tilde{S}$ be as in (14). By convention, if $\mathbf{n}=0$, $A_{\delta}=S$. Let ${\gamma}_{1},...,{\gamma}_{\mathbf{n}}$ be elements of the fundamental group which can be represented by oriented embedded loops around the respective punctures $p_{1},...,p_{\mathbf{n}}$. Those elements are uniquely determined up to conjugacy. Those loops are freely homotopic to the boundary components of $A_{\delta}$.
Let
$\Pi:{\tilde{S}}\to S$
be the natural projection. Consider the complement $\tilde{A}_{\delta}^{c}$ of $\tilde{A}_{\delta}$ inside $\tilde{S}$. The region $\tilde{A}_{\delta}^{c}$ is a disjoint union of connected components, each of which is mapped to an embedded punctured disk inside $S$ under the projection $\Pi$. For each $q\in\{1,...,\mathbf{n}\}$, recall that we fixed a conformal parameterization
$P_{p_{q}}:D(p_{q},1)\to\mathbb{D}^{*}$
and defined $D(p_{q},r)$ in (2). Let us also fix a standard conformal diffeomorphism from $\mathbb{D}^{*}$ to the half-infinite cylinder
$\mathcal{Q}:\quad\mathbb{D}^{*}$$\to\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq 0\}/z\mapsto z+2\pi$
$(\theta,r)$ $\mapsto(\theta,\log\frac{1}{r}).$
For each $q\in\{1,...,\mathbf{n}\}$, consider a connected component $C_{\delta,q}$ of $\tilde{A}_{\delta}^{c}$ left invariant by the cyclic subgroup $\langle\gamma_{q}\rangle$ (this component is uniquely defined). By lifting the conformal parameterization $\mathcal{Q}\circ P_{p_{q}}$ to $C_{\delta,q}$, the closure of the connected component $C_{\delta,q}$ is conformal to
$\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}.$
In this conformal parametrization, $\gamma_{q}$ acts by translation $z\mapsto z+2\pi.$ Note that
$\tilde{A}_{\delta}^{c}=\bigcup_{q=1}^{\mathbf{n}}\Pi^{-1}(\Pi(C_{\delta,q})).$
Lemma 3.2.
Let $\delta\in(0,1]$. Consider a smooth ${\pi_{1}(S)}$-equivariant map $\varphi_{0}:\tilde{A}_{\delta}\to\mathbb{S}_{V}$ with respect to $\rho$, such that for some $\bar{L}>0$, for each $q\in\{1,...,\mathbf{n}\}$,
$\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}\leq\bar{L}.$
Then there is a constant $C=C(\bar{L})>0$ such that the following holds:
  1. (1)
    if $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ for all $p\in\mathrm{Punc}_{S}$, then there is a map $\hat{\varphi}\in\mathscr{H}_{S,\rho}$ such that
    $\operatorname{E}(\hat{\varphi}|_{\mathbf{D}_{S}})\leq\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})+C\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}},$
  2. (2)
    if $\dim_{\mathbb{R}}V<\infty$, then there is a map $\hat{\varphi}\in\mathscr{H}_{S,\rho}$ such that
    $\operatorname{E_{\mathrm{ren}}}(\hat{\varphi}|_{\mathbf{D}_{S}})\leq\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})+C\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}.$
Proof.
The lemma is only nontrivial when there are some punctures, so we will assume in this proof that $\mathbf{n}>0$. Let $\delta\in(0,1]$. Consider the given smooth ${\pi_{1}(S)}$-equivariant map $\varphi_{0}:\tilde{A}_{\delta}\to\mathbb{S}_{V}.$ Since $\mathbf{D}_{S}\cap\tilde{A}_{\delta}$ is compact, $\operatorname{E}((\varphi_{0})|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})<\infty.$ We want to extend this map to a map from the whole $\tilde{S}$ with (renormalized) energy controlled in terms of $\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})$ and the boundary integral $\int_{\mathbf{D}_{S}\cap\partial\tilde{A}_{\delta}}|d\varphi_{0}|^{2}$.
Case $(1)$: The assumption that $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ for all $p\in\mathrm{Punc}_{S}$ means the following: for each $j\geq 1$, there is a vector $v_{q,j}\in\mathbb{S}_{V}$ and some length-minimizing geodesic segment $\beta_{j}$ from $v_{q,j}$ to $\rho(\gamma_{q})v_{q,j}$ with length converging to 00 as $j\to\infty$. Consider a small constant
$\alpha\in(0,10^{-10})$
which will be fixed later. After taking a subsequence in $j$, we can assume without loss of generality that
(16) $\operatorname{Length}(\beta_{j})\leq\alpha e^{-j}.$
Furthermore, by changing $v_{q,j}$ to its opposite $-v_{q,j}$, we can assume without loss of generality that for any $j$
(17) $\operatorname{dist}_{\mathbb{S}_{V}}(v_{q,j},v_{q,j+1})\leq\frac{\pi}{2}.$
We can also assume (after taking the opposite of the $v_{q,j}$’s if necessary) that
(18) $\operatorname{dist}_{\mathbb{S}_{V}}(\varphi_{0}((0,\log\frac{1}{\delta})),v_{q,1})\leq\frac{\pi}{2}.$
Fix momentarily $q\in\{1,...,\mathbf{n}\}$ and consider a corresponding component $C_{\delta,q}$ of $\tilde{A}_{\delta}^{c}$ as before the lemma. By a slight abuse of notations, we identify any map $\phi:C_{\delta,q}\to\mathbb{S}_{V}$ (resp. $\phi:\partial C_{\delta,q}\to\mathbb{S}_{V}$) with the corresponding map
$\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}\to\mathbb{S}_{V}\quad(\text{resp.}\quad\partial\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}\to\mathbb{S}_{V})$
induced by the conformal parametrization fixed above. We will first define a new map $\varphi_{1}$ on $\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}$, which we require to be equivariant, to coincide with $\varphi_{0}$ on the boundary $\partial\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}$, and such that the area of the image of $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$ has controlled area. Then, we will use a reparameterization trick to obtain a map $\varphi_{2}$ with controlled energy on $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$.
We start by setting $\varphi_{1}=\varphi_{0}$ on the boundary $\partial\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}=\mathbb{R}\times\{\log\frac{1}{\delta}\}$. By an arbitrarily small perturbation of $\varphi_{0}$, we can assume (without loss of generality for the proof) that the restriction of $\varphi_{1}$ to $\partial\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq 1\}=\mathbb{R}\times\{1\}$ is an immersion. On the square $[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+1]\subset\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}$, we extend $\varphi_{1}$ in such a way that
  • $\varphi_{1}(.,\log\frac{1}{\delta}+1)$ is a parameterization proportional to arclength of the curve $\varphi_{1}(.,\log\frac{1}{\delta})$, in other words
[…]
  • when $s\in[\log\frac{1}{\delta},\log\frac{1}{\delta}+1]$, $\varphi_{1}(.,s)$ is the natural linear interpolation between the two parameterizations $\varphi_{1}(.,\log\frac{1}{\delta})$ and $\varphi_{1}(.,\log\frac{1}{\delta}+1)$, namely if we set $\gamma(.):=\varphi_{1}(.,\log\frac{1}{\delta}+1)$ and view it as a bijection onto its image,
    $\forall(t,s)\in[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+1],\quad\varphi_{1}(t,s)=\gamma\big{[}(s-\log\frac{1}{\delta})t+(1-s+\log\frac{1}{\delta})\gamma^{-1}(\varphi_{1}(t,\log\frac{1}{\delta}))\big{]}.$
By those two properties, plus the assumption in the lemma, we estimate easily that
(19) $\int_{[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+1]}|d\varphi_{1}|^{2}\leq C_{(0)}\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}\leq C_{(0)}\sqrt{\bar{L}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}$
for some constant $C_{(0)}$ depending only on $\bar{L}$. Here the energy is computed using the standard Euclidean metric $g_{\mathrm{Eucl}}$ on $[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+1]$.
Next set
$\varphi_{1}:[0,2\pi]\times\{\log\frac{1}{\delta}+2\}\to\mathbb{S}_{V}$
to be the geodesic $\beta_{2}$ from $\varphi_{1}(0,\log\frac{1}{\delta}+2):=v_{q,2}$ to $\varphi_{1}(2\pi,\log\frac{1}{\delta}+2):=\rho(\gamma_{q})v_{q,2}$, parametrized proportionally to arclength. From the previous paragraph, $\gamma(.):=\varphi_{1}(.,\log\frac{1}{\delta}+1):[0,2\pi]\to\mathbb{S}_{V}$ is a parameterization proportional to arclength of the curve $\varphi_{1}(.,\log\frac{1}{\delta})$, and
[…]
We claim that for some constant $C_{(1)}$ depending only on $\bar{L}$, we can extend $\varphi_{1}$ to the region $[0,2\pi]\times[\log\frac{1}{\delta}+1,\log\frac{1}{\delta}+2]$ with a map whose differential satisfies at all $(x,y)\in[0,2\pi]\times[\log\frac{1}{\delta}+1,\log\frac{1}{\delta}+2]$:
(21) $\begin{split}|\frac{\partial\varphi_{1}}{\partial x}|&\leq C_{(1)}\big{[}(1-(y-\log\frac{1}{\delta}-1))\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+(y-\log\frac{1}{\delta}-1)\operatorname{Length}(\beta_{2})\big{]}\\ |\frac{\partial\varphi_{1}}{\partial y}|&\leq C_{(1)}.\end{split}$
If the length of $\gamma$ is smaller than say ${10^{-10}}$, such an extension can be cosntructed for instance by interpolating linearly between $\gamma$ and $\beta_{2}$ inside the Hilbert space $V$, using (18) and radially projecting the interpolation to the sphere $\mathbb{S}_{V}$. If the length of $\gamma$ is between ${10^{-10}}$ and $\sqrt{2\pi}\sqrt{\bar{L}}$ (which is always an upper bound by (20)), such an extension exists by an argument by contradiction and a compactness argument. This justifies the claim.
For each $j\geq 3$, set
$\varphi_{1}:[0,2\pi]\times\{\log\frac{1}{\delta}+j\}\to\mathbb{S}_{V}$
to be the geodesic $\beta_{j}$ from $\varphi_{1}(0,\log\frac{1}{\delta}+j):=v_{q,j}$ to $\varphi_{1}(2\pi,\log\frac{1}{\delta}+j):=\rho(\gamma_{q})v_{q,j}$, parametrized proportionally to arclength. For any $j\geq 2$ and $t\in[0,2\pi]$, define
$\varphi_{1}:\{t\}\times[\log\frac{1}{\delta}+{j},\log\frac{1}{\delta}+j+1]\to\mathbb{S}_{V}$
to be a geodesic segment of length less than $\pi$ from $\varphi_{1}((t,\log\frac{1}{\delta}+j))$ to $\varphi_{1}((t,\log\frac{1}{\delta}+j+1))$ parametrized proportionally to arclength. By (16) and (17), $\varphi_{1}$ is well-defined and smooth on $[0,2\pi]\times[\log\frac{1}{\delta}+j,\log\frac{1}{\delta}+j+1]$. Moreover, it is elementary to check that for some constant $C_{(2)}$ depending only on $\bar{L}$, at any point $(x,y)\in[0,2\pi]\times[\log\frac{1}{\delta}+j,\log\frac{1}{\delta}+j+1]$:
(22) $\begin{split}|\frac{\partial\varphi_{1}}{\partial x}|&\leq C_{(2)}\big{[}(1-(y-\log\frac{1}{\delta}-j))\operatorname{Length}(\beta_{j})+(y-\log\frac{1}{\delta}-j)\operatorname{Length}(\beta_{j}+1)\big{]}\\ |\frac{\partial\varphi_{1}}{\partial y}|&\leq C_{(2)}.\end{split}$
Now, the map $\varphi_{1}$ is extended to the whole band $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$, but it may not have controlled energy. On the other hand, it has bounded differential with respect to a metric which we construct now. We denote by $g_{\mathrm{Eucl}}$ the standard Euclidean metric on $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$. Let $g_{1}$ be the piecewise smooth Riemannian metric on $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$ defined as follows:
  • on $[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+1]$,
    $g_{1}=\big{(}\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}\big{)}g_{\mathrm{Eucl}},$
  • on $[0,2\pi]\times[\log\frac{1}{\delta}+1,\log\frac{1}{\delta}+2]$,
    $g_{1}=\big{[}(1-(y-\log\frac{1}{\delta}-1))\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+(y-\log\frac{1}{\delta}-1)\operatorname{Length}(\beta_{2})\big{]}^{2}dx^{2}+dy^{2},$
  • for each $j\geq 2$, on $[0,2\pi]\times[\log\frac{1}{\delta}+j,\log\frac{1}{\delta}+j+1]$,
    $g_{1}=\big{[}(1-(y-\log\frac{1}{\delta}-j))\operatorname{Length}(\beta_{j})+(y-\log\frac{1}{\delta}-j)\operatorname{Length}(\beta_{j+1})\big{]}^{2}dx^{2}+dy^{2}.$
Note that this metric induces a metric on the cylinder $\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}/{z\mapsto z+2\pi}$ which is rotationally symmetric. Here is the key property of this metric: thanks to (21) and (22), we have with respect to $g_{1}$:
$|d\varphi_{1}|_{g_{1}}\leq C_{(3)}$
on $[0,2\pi]\times[\log\frac{1}{\delta}+1,\infty)$ for a constant $C_{(3)}$ depending only on $\bar{L}$. In particular,
[…]
where the first term on the right is due to (19) and conformal invariance of the energy. Because of the lengths bound for $\beta_{j}$ (16) and using the definition of $g_{1}$, we have for some constant $C_{(4)}$ depending only on $\bar{L}$,
$\operatorname{Area}([0,2\pi]\times[\log\frac{1}{\delta}+1,\infty),g_{1})\leq C_{(4)}(\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\alpha)$
which implies with the previous inequality that
[…]
In order to get a map with controlled energy with respect to $g_{\mathrm{Eucl}}$ instead of $g_{1}$, we can use a reparameterization trick. Since $g_{1}$ is rotationally symmetric on the cylinder $\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}/{z\mapsto z+2\pi}$, this Riemannian cylinder is conformal to a flat cylinder $\mathcal{C}$ which is either bounded or half-infinite, via a conformal map which is itself equivariant by rotations of the cylinder. Since the metric $g_{1}$ is of the form $f(y)^{2}dx^{2}+dy^{2}$ on $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$ for all $y\in[\log\frac{1}{\delta},\infty)$ large enough, where $|f(y)|\leq 1$ by the length bound (16) for the $\beta_{j}$’s, the flat cylinder $\mathcal{C}$ is necessarily half-infinite and we can take it to be $\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}/{z\mapsto z+2\pi}$ with the standard metric $g_{\mathrm{Eucl}}$. After lifting to the universal covers, we obtain an equivariant, conformal diffeomorphism
:(×$\mathcal{R}:(\mathbb{R}\times[\log\frac{1}{\delta},\infty),g_{\mathrm{Eucl}})\to(\mathbb{R}\times[\log\frac{1}{\delta},\infty),g_{1})$
which restricts to the identity map on the boundary ×$\mathbb{R}\times\{\log\frac{1}{\delta}\}$.
Consider the new map
$\varphi_{2}:=\varphi_{1}\circ\mathcal{R}:[0,2\pi]\times[\log\frac{1}{\delta},\infty)\to\mathbb{S}_{V}.$
By (23) and conformal invariance of the energy, this map satisfies for some constant $C_{(5)}$ depending only on $\bar{L}$:
(24) $\int_{[0,2\pi]\times[\log\frac{1}{\delta},\infty)}|d\varphi_{2}|^{2}\leq C_{(5)}(\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\alpha)$
where the left-hand side is computed with respect to the Euclidean metric $g_{\mathrm{Eucl}}$.
The map $\varphi_{2}$ defines, after imposing equivariance, a map from $\Pi^{-1}(\Pi(C_{\delta,q}))$ to $\mathbb{S}_{V}$. Repeating the whole construction for each $q\in\{1,...,\mathbf{n}\}$, we obtain an equivariant map
$\varphi_{2}:\tilde{A}^{c}_{\delta}\to\mathbb{S}_{V}$
such that
(25) $\begin{split}\int_{\mathbf{D}_{S}\cap\tilde{A}^{c}_{\delta}}|d\varphi_{2}|^{2}&\leq C_{(5)}\sum_{q=1}^{\mathbf{n}}\big{(}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\alpha\big{)}\\ &\leq C_{(5)}\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\mathbf{n}\alpha\end{split}$
where the energy is computed with respect to the Euclidean metric $g_{\mathrm{Eucl}}$. We can glue the maps $\varphi_{0}$ and $\varphi_{2}$ along the boundary of $\tilde{A}_{\delta}$ to obtain our final equivariant map $\hat{\varphi}:\tilde{S}\to\mathbb{S}_{V}.$ It belongs to $\mathscr{H}_{S,\rho}$ after a smoothing, and its energy satisfies by (25):
$\operatorname{E}(\hat{\varphi}|_{\mathbf{D}_{S}})\leq\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})+C_{(5)}\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\mathbf{n}\alpha$
for a constant $C_{(5)}$ depending only on $\bar{L}$. The positive constant $\alpha>0$ could have been taken arbitrarily small. Either $\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}>0$ and we can chose $\mathbf{n}\alpha$ to be smaller. Or $\sum_{q=1}^{\mathbf{n}}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}=0$, which means that $\varphi_{0}$ is constant on connected components of $\partial\tilde{A}_{\delta}$, so we could have taken the trivial extension instead of $\hat{\varphi}$ above. The trivial extension has same energy as $\varphi_{0}$. In both cases, we conclude Case (1).
Case $(2)$: Let $\delta\in(0,1]$. We use the same notations as above. We also introduce the following notation. Given $q\in\{1,...,\mathbf{n}\}$, we have identified $C_{\delta,q}$ with $\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}.$ Set
$\mathbf{r}:\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}\to[0,\infty)$
$\mathbf{r}:z\mapsto\exp(-\operatorname{Im}(z)).$
In particular, $\mathbf{r}(z)=\delta$ for $z\in\partial\{z\in\mathbb{C};\quad\operatorname{Im}(z)\geq\log\frac{1}{\delta}\}$.
Fix $q\in\{1,...,\mathbf{n}\}$. By the assumption that $\dim_{\mathbb{R}}V<\infty$ and by compactness, there is a point $v_{q}\in\mathbb{S}_{V}$ and some length-minimizing geodesic segment parameterized by arclength
$\beta:[0,2\pi]\to\mathbb{S}_{V}$
such that $\beta(0)=v_{q}$ and $\beta(1)=\rho(\gamma_{q})v_{q}$, and
$\operatorname{Length}(\beta)=\operatorname{\boldsymbol{\lambda}}_{\rho}(q).$
By changing $v_{q}$ to its opposite $-v_{q}$, we can assume without loss of generality that
(26) $\operatorname{dist}_{\mathbb{S}_{V}}(\varphi_{0}((0,\log\frac{1}{\delta})),v_{q})\leq\frac{\pi}{2}.$
Next, we want to define an extension of $\varphi_{0}$ with controlled energy. We first extend $\varphi_{0}$ to a map $\varphi_{0}^{\prime}$ on $[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+2]$ such that
$\varphi_{0}^{\prime}(.,\log\frac{1}{\delta}+2)=\beta.$
Similar to the proof of Case (1), we choose $\varphi_{0}^{\prime}$ to have controlled energy and area on $[0,2\pi]\times[\log\frac{1}{\delta},\log\frac{1}{\delta}+2]$, and the reparameterization trick enables us to construct another extension $\varphi_{1}:[0,2\pi]\times[\log\frac{1}{\delta},h]\to\mathbb{S}_{V}$ for some $h>\log\frac{1}{\delta}$, such that $\varphi_{1}(.,h)=\varphi_{0}^{\prime}(.,\log\frac{1}{\delta}+2)$ and for a constant $C_{(6)}$ depending only on $\bar{L}$:
(27) $\begin{split}\int_{[0,2\pi]\times[\log\frac{1}{\delta},h]}|d\varphi_{1}|^{2}&\leq C_{(6)}\big{(}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}+\operatorname{\boldsymbol{\lambda}}_{\rho}(q)\big{)}\\ &\leq C_{(6)}\big{(}1+\sqrt{2\pi})\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}\end{split}$
where we used that $\operatorname{\boldsymbol{\lambda}}_{\rho}(q)=\operatorname{Length}(\beta)\leq\sqrt{2\pi}\sqrt{\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}}$ by Cauchy-Schwarz. Then we extend $\varphi_{1}$ to $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$ by setting for all $(t,s)\in[0,2\pi]\times[h,\infty)$,
$\varphi_{1}(t,s)=\beta(t).$
The “renormalized” energy of $\varphi_{1}$ on $[0,1]\times[h,\infty)$ satisfies
$\lim_{r\to 0}\big{[}\big{(}\frac{1}{2}\int_{[0,2\pi]\times[h,\infty)\setminus\mathbf{r}^{-1}((0,r))}|d\varphi_{1}|^{2}\big{)}-\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p_{q})^{2}}{4\pi}\log\frac{1}{r}\big{]}=-\frac{\operatorname{\boldsymbol{\lambda}}_{\rho}(p_{q})^{2}}{4\pi}h\leq 0.$
The “renormalized” energy of $\varphi_{1}$ on $[0,2\pi]\times[\log\frac{1}{\delta},\infty)$ is thus at most $\int_{[0,2\pi]\times[\log\frac{1}{\delta},h]}|d\varphi_{1}|^{2}$, which is controlled by (27). Repeating this construction for every cusp, defining $\varphi_{1}$ on the whole universal cover $\tilde{S}$ by equivariance, and adding the previous estimates together, we conclude Case (2).
The following corollary contains a practical characterization of pairs $(S,\rho)$ with finite energy, and gives an a priori bound on $\operatorname{E_{\mathrm{ren}}}(S,\rho)$.
Corollary 3.3.
  1. (1)
    $\operatorname{E}(S,\rho)<\infty$ if and only if $\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ for all $p\in\mathrm{Punc}_{S}$.
  2. (2)
    If $\dim_{\mathbb{R}}V<\infty$, the renormalized energy $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ is finite and bounded from above by a constant $C_{S}$ depending only on $S$.
Proof.
Fix a triangulation of $S$ and any $\delta\in(0,1]$. By basic topological arguments, there is always a smooth ${\pi_{1}(S)}$-equivariant map $\varphi_{0}:\tilde{A}_{\delta}\to\mathbb{S}_{V}$. Moreover, since the diameter of $\mathbb{S}_{V}$ is $\pi$, by straightening the map on the 1-skeleton of the triangulation then extending the map by hand to the 2-skeleton, $\varphi_{0}$ can be chosen so that $\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}})\leq C^{\prime}_{S}$ and $\int_{\mathbf{D}_{S}\cap\partial{C_{\delta,q}}_{\delta}}|d\varphi_{0}|^{2}<C^{\prime}_{S}$ for all $q\in\{1,...,\mathbf{n}\}$, for some $C^{\prime}_{S}>0$ independent of $\rho$ or the dimension of $V$. Here $C_{\delta,q}$ is defined before Lemma 3.2.
(1): The implication $\forall p\in\mathrm{Punc}_{S},\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0\Rightarrow\operatorname{E}(S,\rho)<\infty$ is an easy consequence of Lemma 3.2 (1) and the existence of $\varphi_{0}$ explained above. Next, the direction $\operatorname{E}(S,\rho)<\infty\Rightarrow\forall p\in\mathrm{Punc}_{S},\operatorname{\boldsymbol{\lambda}}_{\rho}(p)=0$ follows from Fubini’s theorem and was already showed in the proof of Lemma 1.4.
(2): This follows from Lemma 3.2 (2) and the existence of $\varphi_{0}$ explained above.
Corollary 3.4.
  1. (1)
    If $\rho:{\pi_{1}(S)}\to\operatorname{End}(V)$, $\nu:{\pi_{1}(S)}\to\operatorname{End}(K)$ are unitary representations such that $\rho\prec\nu$, then
    $\operatorname{E}(S,\bigoplus^{\infty}\nu)\leq\operatorname{E}(S,\rho).$
  2. (2)
    If moreover $\rho$ is irreducible, $\nu=\bigoplus_{m=1}^{\infty}\nu_{m}$, then for any $\epsilon^{\prime\prime}>0$, there is $m_{0}\geq 1$ such that
    $\operatorname{E_{\mathrm{ren}}}(S,\nu_{m_{0}})\leq\operatorname{E}(S,\rho)+\epsilon^{\prime\prime}.$
Proof.
(1): Suppose $\operatorname{E}(S,\rho)<\infty$, otherwise there is nothing to do. Consider a map $\varphi_{0}\in\mathscr{H}_{S,\rho}$ with finite energy on the fundamental domain $\mathbf{D}_{S}$. By Fubini’s theorem as in the proof of Lemma 1.4, for any $\epsilon>0$, there is an arbitrarily small $\delta>0$ such that for any $q\in\{1,...,\mathbf{n}\}$,
$\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}|^{2}\leq\epsilon$
where $C_{\delta,q}$ is defined before Lemma 3.2. By Proposition 3.1 (1), we get a smooth map $\varphi_{0}^{\prime}:\tilde{A}_{\delta}\to\mathbb{S}_{\bigoplus^{\infty}K}$, which is $\pi_{1}(S)$-equivariant with respect to $\bigoplus^{\infty}\nu$, such that
$|\operatorname{E}(\varphi_{0}^{\prime}|_{\mathbf{D}_{S}})-\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}})|<\epsilon\quad\text{and}\quad\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}^{\prime}|^{2}\leq 2\epsilon\quad\text{for any $q\in\{1,...,\mathbf{n}\}$}.$
Applying Lemma 3.2 (1) and letting $\epsilon\to 0$, we obtain $\operatorname{E}(S,\bigoplus^{\infty}\nu)\leq\operatorname{E}(S,\rho)$.
(2): Suppose moreover that $\rho$ is irreducible, and $\nu=\bigoplus_{m=1}^{\infty}\nu_{m}$ with $\nu_{m}:\pi_{1}(S)\to\operatorname{End}(K_{m})$. Using Proposition 3.1 (2), for any $\epsilon>0$, we get an $m_{0}=m_{0}(\epsilon)\geq 1$ and a smooth map $\varphi_{0}^{\prime\prime}:\tilde{A}_{\delta}\to\mathbb{S}_{K_{m_{0}}}$, which is $\pi_{1}(S)$-equivariant with respect to $\nu_{m_{0}}$, such that
$|\operatorname{E}(\varphi_{0}^{\prime\prime}|_{\mathbf{D}_{S}})-\operatorname{E}(\varphi_{0}|_{\mathbf{D}_{S}})|<\epsilon\quad\text{and}\quad\int_{\mathbf{D}_{S}\cap\partial C_{\delta,q}}|d\varphi_{0}^{\prime\prime}|^{2}\leq 2\epsilon\quad\text{for any $q\in\{1,...,\mathbf{n}\}$}.$
Applying Lemma 3.2 (2) and choosing $\epsilon$ small enough, we have $\operatorname{E_{\mathrm{ren}}}(S,\nu_{m_{0}})\leq\operatorname{E}(S,\rho)+\epsilon^{\prime\prime}$ for any $\epsilon^{\prime\prime}>0$ fixed. ∎
4. From strong convergence to convergence of harmonic maps
We prove the main result, Theorem 0.4, in this section. Let us recall our assumptions and notations. Let $S$ be a punctured Riemann surface, with genus $\mathbf{g}$ and $\mathbf{n}\geq 0$ punctures. Suppose that $S$ has negative Euler characteristic $\chi(S)<0$. Let
$\rho_{j}:\pi_{1}(S)\to U(N_{j})$
be a sequence of unitary representations.
4.1. Convergence of the renormalized energy
In this subsection, we show that the first part of Theorem 0.4: strong convergence of a sequence of representations (see Subsection 6.2 in the Appendix) implies convergence of the renormalized energy.
Theorem 4.1.
If $\rho_{j}$ strongly converges, then
$\lim_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)=\frac{\pi}{4}|\chi(S)|.$
Note that $\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$ is exactly $\frac{1}{8}\operatorname{Area}(S,g_{\mathrm{hyp}})$, where $g_{\mathrm{hyp}}$ is the unique complete, finite area, hyperbolic metric on $S$ compatible with its conformal structure. The theorem above will follow from Proposition 4.2 and Proposition 4.4 below.
Proposition 4.2.
Under the previous assumptions,
$\limsup_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})\leq\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2).$
Proof.
Arguing towards a contradiction, let us assume that for some $\epsilon^{\prime\prime}>0$, there is a sequence $\{j_{m}\}_{m\geq 1}$ converging to $\infty$ such that
(28) $\inf_{m\geq 1}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j_{m}})>\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)+\epsilon^{\prime\prime}.$
Set
$\nu:=\bigoplus_{m\geq 1}\rho_{j_{m}}.$
Let $\mathscr{P}:\tilde{S}\to\mathbb{S}_{2}(\partial\tilde{S})$ and $\underline{\rho}_{B}$ be, respectively, the map in $\mathscr{H}_{S,\underline{\rho}_{B}}$ and boundary representation defined in (8) and (7). Note that since from Lemma 1.13 (1), $\underline{\rho}_{B}\sim\lambda_{\pi_{1}(S)}$, we get
$\|\underline{\rho}_{B}(f)\|=\|\lambda_{\pi_{1}(S)}(f)\|\quad\text{for all $f\in\mathbb{C}[{\pi_{1}(S)}]$}$
by Theorem 6.2. Since $\rho_{j}$ strongly converges by assumption,
$\lim_{j\to\infty}\|\rho_{j}(f)\|=\|\lambda_{\pi_{1}(S)}(f)\|\quad\text{for all$f\in\mathbb{C}[{\pi_{1}(S)}]$}.$
Hence, we clearly have
$\|\underline{\rho}_{B}(f)\|\leq\|\nu(f)\|\quad\text{for all $f\in\mathbb{C}[{\pi_{1}(S)}]$}$
which means according to Theorem 6.2 that $\underline{\rho}_{B}\prec\nu$. Next, we know from Lemma 1.13 that $\underline{\rho}_{B}$ is irreducible and $\operatorname{E}(\mathscr{P}|_{\mathbf{D}_{S}})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$. We can then apply Corollary 3.4 (2) to $\rho:=\underline{\rho}_{B}$, $\nu:=\bigoplus_{m\geq 1}\rho_{j_{m}}$, and conclude that for some $m_{0}\geq 1$,
$\operatorname{E_{\mathrm{ren}}}(S,\rho_{j_{m_{0}}})\leq\operatorname{E}(S,\underline{\rho}_{B})+\epsilon^{\prime\prime}\leq\operatorname{E}(\mathscr{P}|_{\mathbf{D}_{S}})+\epsilon^{\prime\prime}=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)+\epsilon^{\prime\prime}.$
That contradicts (28), as wanted. ∎
We observe an easy consequence of the strong convergence of $\rho_{j}$ (see (3) for the definition of $\operatorname{\boldsymbol{\lambda}}_{\rho_{j}}(p)$):
Lemma 4.3.
If $\rho_{j}$ strongly converges, for any puncture $p\in\mathrm{Punc}_{S}$, we have
$\lim_{j\to\infty}\operatorname{\boldsymbol{\lambda}}_{\rho_{j}}(p)=0.$
Proof.
Indeed, consider any element $g\in\pi_{1}(S)$ in the conjugacy class determined by small embedded oriented loops around $p$ in $S$. For any $\varepsilon>0$, there is a function $F_{\varepsilon}\in\ell^{2}(\pi_{1}(S))$ of unit $L^{2}$-norm such that $\|\lambda_{\pi_{1}(S)}(g)F_{\varepsilon}-F_{\varepsilon}\|\leq\varepsilon$ where $\lambda_{\pi_{1}(S)}$ is the regular representation (e.g. take $F_{\varepsilon}$ to be constant on its support and supported on $\{g,g^{2},....,g^{N_{\varepsilon}}\}$ for some large integer $N_{\varepsilon}$). This implies in terms of operator norm: $\|\lambda_{\pi_{1}(S)}(g)+\lambda_{\pi_{1}(S)}(e)\|=2$ where $e\in\pi_{1}(S)$ is the identity element. By strong convergence, we have $\lim_{j\to\infty}\|\rho_{j}(g)+\rho_{j}(e)\|=2$. But this means by Cauchy-Schwarz that for any $\varepsilon>0$, for all large $j$, there are unit norm vectors $v_{j}$ such that $\|\rho_{j}(g)v_{j}-v_{j}\|\leq\varepsilon$, which is exactly the lemma.
The next proposition treats the opposite inequality. Its proof is the key place where the concept of strong convergence for representations is combined with harmonic maps. There, we construct a limit of harmonic maps, equivariant with respect to a limit representation, which we show is weakly equivalent to the regular representation.
Proposition 4.4.
Under the above assumptions,
$\liminf_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})\geq\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2).$
Proof.
For each $j\geq 1$, let $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ be a harmonic representative of $(S,\rho_{j})$: it is a harmonic map in $\mathscr{H}_{S,\rho_{j}}$ given by Theorem 1.7, so that
$\operatorname{E_{\mathrm{ren}}}(\psi_{j}|_{\mathbf{D}_{S}})=\operatorname{E_{\mathrm{ren}}}(S,\rho_{j}).$
Due to Lemma 1.5, on every compact subset of $\mathbf{D}_{S}$, the energy of $\psi_{j}$ is uniformly bounded since $\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})$ is uniformly bounded thanks to Corollary 3.3 (2).
By standard harmonic map theory [SU82], after picking a subsequence and using the ${\pi_{1}(S)}$-equivariance of the map $\psi_{j}$, a priori there are finitely many points $y_{1},...,y_{k}\in\mathbf{D}_{S}$ where the energy might concentrate as $j\to\infty$. Using the $\epsilon$-regularity theorem, Theorem 1.8, we can replace $\psi_{j}$ in a small neighborhood of the ${\pi_{1}(S)}$-orbit of $\bigcup_{j=1}^{k}y_{j}$ equivariantly by a map with renormalized energy strictly smaller than $\psi_{j}$ for $j$ large. But since $\psi_{j}$ is energy minimizing among all smooth equivariant maps, we have a contradiction (this is the usual replacement argument appearing in [SU82, proof of Theorem 5.1] for instance). Thus, the energy of $\psi_{j}$ does not concentrate around any point as $j\to\infty$. Next, by the $\epsilon$-regularity theorem again, Theorem 1.8, on each compact subset of $\tilde{S}$, the maps $\psi_{j}$ are uniformly bounded in the $C^{m}$-topology for any $m\geq 0$.
In order to construct a limit, let us view all the spaces $\mathbb{C}^{N_{j}}$ as embedded as linear subspaces of one common Hilbert space $H$ with unit sphere $\mathbb{S}_{H}$. Choose a dense countable sequence of points $\{p_{t}\}_{t\geq 0}$ in $\tilde{S}$. Then, since the unitary group of $H$ acts transitively on (complex) orthonormal bases of $H$, there is a sequence of unitary transformations $F_{j}$ of $H$ such that (after taking a subsequence in $j$ if necessary) for any given $t\geq 0$, the sequence $F_{j}\circ\psi_{j}(p_{t})$ converges to a point in $\mathbb{S}_{H}$ as $j\to\infty$. Since the maps $\psi_{j}$ are uniformly Lipschitz on compact subsets, by an Arzelà-Ascoli argument, $F_{j}\circ\psi_{j}$ converges in the $C^{0}$-topology on compact subsets to a limit locally Lipschitz map
$\psi_{\infty}:\tilde{S}\to\mathbb{S}_{H}.$
[…]
We now bound the energy of the limit map $\psi_{\infty}$. By the lower semicontinuity of the energy under convergence of harmonic maps, Lemma 1.5 and Lemma 4.3:
$\begin{split}\operatorname{E}(\psi_{\infty}|_{\mathbf{D}_{S}})&=\lim_{\delta\to 0}\operatorname{E}(\psi_{\infty}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})\\ &\leq\lim_{\delta\to 0}\liminf_{j\to\infty}\operatorname{E}(\psi_{j}|_{\mathbf{D}_{S}\cap\tilde{A}_{\delta}})\\ &=\lim_{\delta\to 0}\liminf_{j\to\infty}[\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})+\sum_{p\in\mathrm{Punc}_{S}}\frac{\operatorname{\boldsymbol{\lambda}}_{\rho_{j}}(p)^{2}}{4\pi}\log\frac{1}{\delta}]\\ &=\liminf_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j}).\end{split}$
Hence,
(29) $\operatorname{E}(S,\rho^{\prime})\leq\operatorname{E}(\psi_{\infty}|_{\mathbf{D}_{S}})\leq\liminf_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j}).$
Next, let us explain how the limit map $\psi_{\infty}$ determines a unitary representation $\rho^{\prime}:{\pi_{1}(S)}\to\operatorname{End}(H).$ For the sake of simplicity, we write “converge” and “$\lim_{j\to\infty}$” instead of “subsequentially converge”.
By reducing the Hilbert space $H$, we will assume without loss of generality that no closed linear (strict) subspace of $H$ contains $\psi_{\infty}(\tilde{S})$. In other words, there is a sequence of points $z_{1},z_{2},...,z_{i},...\in\tilde{S}$, such that the vectors $\psi_{\infty}(z_{1}),\psi_{\infty}(z_{2}),...\in\psi_{\infty}(\tilde{S})$ are linearly independent, and after applying the Gram-Schmidt procedure, we get a Hilbert orthonormal basis $e_{1}=\psi_{\infty}(z_{1}),e_{2},...,e_{i},...$ of $H$. Note that $F_{j}\circ\psi_{j}(z_{i})$ converges to $\psi_{\infty}(z_{i})$. Thus, for any finite set of complex numbers $a_{j}$, $\sum_{i}a_{i}F_{j}\circ\psi_{j}(z_{i})$ converges to $\sum_{i}a_{i}\psi_{\infty}(z_{i})$. Next, for any $g\in{\pi_{1}(S)}$, and any $z_{i}$, set
(30) $\rho^{\prime}(g)\psi_{\infty}(z_{i}):=\psi_{\infty}(g.z_{i})=\lim_{j\to\infty}(F_{j}\circ\rho_{j}(g)\circ F_{j}^{-1})\circ F_{j}\circ\psi_{j}(z_{i}).$
Since $e_{i}$ is a finite linear combination of $\psi_{\infty}(z_{1}),\psi_{\infty}(z_{2}),...,\psi_{\infty}(z_{i})$, the above formula defines by linearity a linear transformation $\rho^{\prime}(g):H\to H$ which is easily checked to be unitary, since $\rho_{N}(g)$ is unitary. Moreover, by the formula above again, we clearly have $\rho^{\prime}(gg^{\prime})=\rho^{\prime}(g)\rho^{\prime}(g^{\prime})$ since the analogue property holds for $F_{j}\circ\rho_{j}(.)\circ F_{j}^{-1}$. This finishes the definition of the unitary representation $\rho^{\prime}$ and the map $\psi_{\infty}$, which by construction is ${\pi_{1}(S)}$-equivariant with respect to $\rho^{\prime}$, namely $\psi_{\infty}$ is an element of $\mathscr{H}_{S,\rho^{\prime}}$:
(31) $\rho^{\prime}:{\pi_{1}(S)}\to\operatorname{End}(H),\quad\psi_{\infty}:\tilde{S}\to\mathbb{S}_{H}.$
We claim that for all $f\in\ell^{1}({\pi_{1}(S)},\mathbb{C})$,
(32) $\|\rho^{\prime}(f)\|\leq\limsup_{j}\|\rho_{j}(f)\|.$
Indeed, by (30) and using the previous notations, for any finite linear combination $f=\sum_{k}\alpha_{k}g_{k}\in\mathbb{C}[{\pi_{1}(S)}]$, and any finite linear combination $\sum_{i}a_{i}\psi_{\infty}(z_{i})$,
$\|\rho^{\prime}(f)\big{(}\sum_{i}a_{i}\psi_{\infty}(z_{i})\big{)}\|$ $=\|\lim_{j\to\infty}(F_{j}\circ\rho_{j}(f)\circ F_{j}^{-1})\big{(}\sum_{i}a_{i}F_{j}\circ\psi_{j}(z_{i})\big{)}\|$
$=\lim_{j\to\infty}\|(F_{j}\circ\rho_{j}(f)\circ F_{j}^{-1})\big{(}\sum_{i}a_{i}F_{j}\circ\psi_{j}(z_{i})\big{)}\|$
$\leq\big{(}\limsup_{j\to\infty}\|\rho_{j}(f)\|\big{)}\|\sum_{i}a_{i}\psi_{\infty}(z_{i})\|.$
That is enough to show the claim (32). Since by assumption $\rho_{j}$ converges strongly, we get from (32) that for all $f\in\ell^{1}({\pi_{1}(S)},\mathbb{C})$:
$\|\rho^{\prime}(f)\|\leq\|\lambda_{\pi_{1}(S)}(f)\|.$
Thus by Theorem 6.2, $\rho^{\prime}\prec\lambda_{\pi_{1}(S)}$ and so by $C^{*}$-simplicity of ${\pi_{1}(S)}$ [Pow75] [dLH07],
(33) $\rho^{\prime}\sim\lambda_{\pi_{1}(S)}.$
By (33), Lemma 1.12, Theorem 1.11 and Corollary 3.4 (1) applied to $\rho:=\rho^{\prime}$ and $\nu:=\lambda_{\pi_{1}(S)}$, we find that
$\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)=\operatorname{E}(S,\bigoplus^{\infty}\lambda_{\pi_{1}(S)})\leq\operatorname{E}(S,\rho^{\prime}).$
Combined with (29), this finishes the proof. For future use, we record the following consequence:
(34) $\operatorname{E}(S,\rho^{\prime})=\operatorname{E}(\psi_{\infty}|_{\mathbf{D}_{S}})=\liminf_{j\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2).$
Remark 4.5.
M. Magee pointed out that, when $\operatorname{E}(S,\rho_{j})$ is finite, Proposition 4.4 alternatively follows from a generalization of the resolvent method of [HM23] applied to flat $\mathbb{C}^{N}$-bundles over hyperbolic surfaces [Zar22] [Hid23]. Our proof of Proposition 4.4 is different and has the advantage of producing a limit harmonic map, which will be crucial in the proof of the main part of Theorem 0.4 in the next subsection.
4.2. Convergence of the pullback metric
To finish the proof of the main theorem, it remains to identify the limit of the pullback metric. For each $N$, let $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ be a harmonic representative of $(S,\rho_{j})$. Let $g_{\mathrm{Eucl}}$ be the standard Euclidean metric on $\mathbb{S}^{2N_{j}-1}$ and let $g_{\mathrm{hyp}}$ be the unique complete, conformal, hyperbolic metric on the Riemannian surface $S$. By equivariance of $\psi_{j}$, the pullback metric $(\psi_{j})^{*}g_{\mathrm{Eucl}}$ descends to a metric on $S$ still denoted by $(\psi_{j})^{*}g_{\mathrm{Eucl}}$.
Theorem 4.6.
If $\rho_{j}$ strongly converges, then the harmonic representatives $\psi_{j}$ satisfy
$\lim_{j\to\infty}(\psi_{j})^{*}g_{\mathrm{Eucl}}=\frac{1}{8}g_{\mathrm{hyp}}$
in the $C^{\infty}$ topology on compact subsets of $S$.
Proof.
Let $\rho^{\prime}:{\pi_{1}(S)}\to\operatorname{End}(H)$, $\psi_{\infty}:\tilde{S}\to\mathbb{S}_{H}$ be the limit unitary representation and limit map constructed in (31) in the proof of Proposition 4.4. The map $\psi_{\infty}$ is the limit of a subsequence of $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$, after composing with unitary transformations $F_{j}$. The convergence is smooth on any compact subset of $\tilde{S}$. By (33), we have $\rho^{\prime}\sim\lambda_{\pi_{1}(S)}$.
By (34), $\operatorname{E}(\psi_{\infty}|_{\mathbf{D}_{S}})=\operatorname{E}(S,\rho^{\prime})=\frac{\pi}{4}(2\mathbf{g}+\mathbf{n}-2)$. By the equality case of Corollary 2.4,
(35) $(\psi_{\infty})^{*}g_{H}=\frac{1}{8}g_{\mathrm{hyp}}$
where $g_{\mathrm{hyp}}$ is the unique hyperbolic metric on $S$ compatible with its conformal structure.
By the way $\psi_{j}$ converges to $\psi_{\infty}$, for any fixed compact subset $G$ of $\tilde{S}$, we have for any $m\geq 0$:
(36) $\lim_{j\to\infty}\|\psi_{j}^{*}g_{\mathrm{Eucl}}-\frac{1}{8}g_{\mathrm{hyp}}\|_{C^{m}(G)}=0.$
This implies the desired conclusion that $\psi_{j}^{*}g_{\mathrm{Eucl}}$, viewed as a Riemannian metric on $S$, converges to $\frac{1}{8}g_{\mathrm{hyp}}$ on compact subsets of $S$ in the $C^{\infty}$ topology.
5. Applications
5.1. Limit of the $N$th average renormalized energy
Let $S$ be a punctured Riemann surface and suppose that $\pi_{1}(S)$ is a free group $F_{k}$ of rank $k\geq 2$, namely $S$ has at least one puncture. Its Euler characteristic is
$\chi(S)=1-k.$
Let $g_{\mathrm{hyp}}$ be its hyperbolic metric. Unitary representations of $F_{k}$ into $U(N)$ are in one-to-one correspondence with $k$-tuple $(u_{1},...,u_{k})\in U(N)^{k}$. For $N\geq 1$, set
$\mathbb{E}_{N}(S):=\text{ average of $\operatorname{E_{\mathrm{ren}}}(S,\rho)$over all unitary representations $\rho:\pi_{1}(S)\to U(N)$}$
where the average is taken with respect to the Haar measure on $U(N)^{k}$. This measure is preserved by the action of the outer automorphism group $\text{Out}(F_{k})$ (see [CMP19, Lemma 1.5]), which means that $\mathbb{E}_{N}(S)$ does not depend on the identification $\pi_{1}(S)\approx F_{k}$, and in fact induces a function on moduli space of Riemann surfaces. When $\pi_{1}(S)$ is not free (a case we will not consider here), the definition of $\mathbb{E}_{N}(S)$ can naturally be generalized by taking the average of $\operatorname{E_{\mathrm{ren}}}(S,\rho)$ on the moduli space of unitary representations with respect to the Atiyah-Bott-Goldman symplectic volume form [AB83] [Gol84].
In the next statement, by “random sequence of unitary representations $\tau_{N}:\pi_{1}(S)\to U(N)$”, we mean a sequence sampled on $\Pi_{N\geq 1}U(N)^{k}$ from the standard product Haar probability measure. Theorem 0.1 is a corollary of the following theorem:
Theorem 5.1.
Let $S$ be a punctured Riemannian surface with $\pi_{1}(S)$ isomorphic to a free group of rank $k\geq 2$. Consider a random sequence of unitary representations $\tau_{N}:\pi_{1}(S)\to U(N)$. For each $N\geq 1$, let $\psi_{N}:\tilde{S}\to(\mathbb{S}^{2N-1},g_{\mathrm{Eucl}})$ be a harmonic representative of $(S,\tau_{N})$. Then
$\lim_{N\to\infty}\operatorname{E_{\mathrm{ren}}}(S,\tau_{N})=\frac{\pi}{4}(k-1)\text{ almost surely}$
and $(\psi_{N})^{*}g_{\mathrm{Eucl}}$ converges to $\frac{1}{8}g_{\mathrm{hyp}}$ almost surely in the $C^{\infty}$-topology on compact sets of $S$. In particular,
$\lim_{N\to\infty}\mathbb{E}_{N}(S)=\frac{\pi}{4}(k-1).$
Proof.
If clearly suffices to prove the first part of the theorem because by Lemma 3.3, $\operatorname{E_{\mathrm{ren}}}(S,\rho)\leq C_{S}$ for a constant depending only on $S$. By [CM14], a random sequence $\tau_{N}:\pi_{1}(S)\to U(N)$ strongly converges almost surely (see Theorem 6.4 in the Appendix). The statement then follows from Theorem 0.4. ∎
5.2. Special immersions of surfaces into $\mathbb{R}^{n}$
Theorem 5.2.
Let $\epsilon>0$. For any closed hyperbolic surface $(\Sigma,g_{\mathrm{hyp}})$, there exists a finite degree covering $(\Sigma^{\prime},g_{\mathrm{hyp}})$ and a harmonic immersion into a Euclidean unit sphere
$\psi:(\Sigma^{\prime},g_{\mathrm{hyp}})\to(\mathbb{S}^{n_{\epsilon}},g_{\mathrm{Eucl}})$
such that $\psi^{*}g_{\mathrm{Eucl}}$ is $\epsilon$-close to $\frac{1}{8}g_{\mathrm{hyp}}$ in the $C^{2}$-topology.
Proof.
Set $S$ be the Riemann surface corresponding to $(\Sigma,g_{\mathrm{hyp}})$. By [LMH25] (see Theorem 6.7 in the Appendix), there is a sequence of unitary representations
$\rho_{j}:\pi_{1}(S)\to U(N_{j})$
which strongly converges, and such that $\rho_{j}(\pi_{1}(S))$ is a finite subgroup of $U(N_{j})$. The kernel $\ker(\rho_{j})$ is thus a finite index subgroup of $\pi_{1}(S)$. This implies that, if $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ is a harmonic representative of $(S,\rho_{j})$, the map $\psi_{j}$ factors as follows: there is a finite covering $S^{\prime}$ of $S$ with fundamental group $\ker(\rho_{j})$, and a map
$\psi^{\prime}_{j}:S^{\prime}\to\mathbb{S}^{2N_{j}-1}$
such that
$\psi_{j}=\psi^{\prime}_{j}\circ\Pi$
where $\Pi:\tilde{S}\to S^{\prime}$ is the natural projection. This $\psi^{\prime}_{j}$ is a harmonic map and by Theorem 0.4, the pullback metric $(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}$ converges to $\frac{1}{8}g_{\mathrm{hyp}}$ and in particular $\psi^{\prime}_{j}$ is an immersion for large $j$. The theorem is proved.
Remark 5.3 (Meaning of $\frac{1}{8}$).
For $\epsilon>0$ small, the harmonic map $\psi$ in Theorem 5.2 is given by $n_{\epsilon}+1$ coordinates functions $(\psi)^{i}:(\Sigma^{\prime},g_{\mathrm{hyp}})\to\mathbb{R}$ which are almost Laplace eigenfunctions with eigenvalue $2$. Indeed, the image $\psi(\Sigma^{\prime})$ is almost a minimal surface in $\mathbb{S}^{n_{\epsilon}}$, and coordinate functions on 2-dimensional minimal surfaces in Euclidean unit spheres are Laplace eigenfunctions with eigenvalue $2$ by [Tak66]. But $2$ is exactly the bottom of the Laplace operator on the rescaled hyperbolic plane $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$. Optimistically, it could be that Theorem 5.2 holds (resp. cannot hold) if the factor $\frac{1}{8}$ is replaced by any larger (resp. smaller) constant.
5.3. Almost hyperbolic minimal surfaces in spheres
Consider a Riemann surface $T$ and
$\varphi:T\to(M,g_{M})$
a branched minimal immersion of $T$ into some Riemannian manifold $(M,g_{M})$. Then the pullback metric $\varphi^{*}g_{M}$ is a nondegenerate Riemannian metric everywhere on $T$ except at a discrete set of points. It induces a well-defined metric on $T$, and its Gaussian curvature, denoted by $K_{\varphi^{*}g_{M}}$, is well-defined almost everywhere. The map $\varphi$ is said to be simple if there is no nontrivial branched covering map $f:T\to\hat{T}$ and there is no map $\hat{\varphi}:\hat{T}\to M$ so that $\varphi=\hat{\varphi}\circ f$.
Theorem 5.4.
For any $j\geq 1$, there are a closed Riemann surface $T_{j}$ and a simple, branched, minimal immersion
$\varphi_{j}:T_{j}\to(\mathbb{S}^{n_{j}},g_{\mathrm{Eucl}})$
such that
  1. (1)
    $\lim_{j\to\infty}\frac{1}{\operatorname{Area}(T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}})}\int_{T_{j}}|K_{\varphi_{j}^{*}g_{\mathrm{Eucl}}}+8|=0$,
  2. (2)
    $(T_{j},\varphi_{j}^{*}g_{{\mathrm{Eucl}}})$ Benjamini-Schramm converges to $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$ as $j\to\infty$.
Benjamini-Schramm convergence was originally a notion of convergence for random graphs. It has been recently extended to manifolds, see for instance Abért-Biringer [AB22, Section 1]. In our statement, it means the following: there are regions $\Omega_{j}\subset T_{j}$ such that
$\lim_{j\to\infty}\frac{\mathrm{Area}\bigl{(}\Omega_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}}\bigr{)}}{\mathrm{Area}\bigl{(}T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}}\big{)}}=1$
and for any $R>0$ and $x_{j}\in\Omega_{j}$, the metric $R$-ball in $\bigl{(}T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}}\bigr{)}$ centered at $x_{j}$ converges in the Gromov-Hausdorff topology to an $R$-ball in $\bigl{(}\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}}\bigr{)}$.
Proof of Theorem 5.4.
For (1): Take $S$ to be the round unit 2-sphere with three points removed. Let $\Sigma$ be the underlying topological surface, which is a thrice-punctured sphere. It is well-known that the Teichmüller space of $\Sigma$ is reduced to a single point $\mathscr{T}_{\Sigma}=\{\mu_{0}\}$, and that the mapping class group of $\Sigma$ (i.e. the set of isotopy classes of the group of orientation-preserving homeomorphisms of $\Sigma$) is isomorphic to symmetric group $\mathbf{S}_{3}$ (the permutation group of $\{1,2,3\}$) [FM11, remark after Proposition 10.5, and Proposition 2.3]. Let $g_{\mathrm{hyp}}$ be the complete finite area hyperbolic metric on $\Sigma$, unique up to isometries. The elements of the mapping class group are represented by the “obvious” symmetries of $(\Sigma,g_{\mathrm{hyp}})$.
We have $\pi_{1}(S)\approx F_{2}$. By [BC19], namely Theorem 6.5 in the Appendix, and Lemma 6.6, there exists a sequence of unitary representations $\rho_{j}:\pi_{1}(S)\to U(N_{j})$ which strongly converges, such that $\rho_{j}(\pi_{1}(S))$ is finite, and such that $\rho_{j}$ satisfies $\operatorname{\boldsymbol{\lambda}}_{\rho_{j}}(p)=0$ for each puncture $p$ of $S$ (see (3) for the definition of $\operatorname{\boldsymbol{\lambda}}_{\rho_{j}}(p)$). In particular
(37) $\operatorname{E}(S,\rho_{j})=\operatorname{E_{\mathrm{ren}}}(S,\rho_{j})<\infty$
by Corollary 3.3 (1) and Lemma 1.4.
By Theorem 1.7, there is a harmonic representative $\psi_{j}:\tilde{S}\to\mathbb{S}^{2N_{j}-1}$ of $(S,\rho_{j})$, so thanks to (37), $\operatorname{E}(\psi_{j}|_{\mathbf{D}_{S}})=\operatorname{E}(S,\rho_{j}).$ Because the Teichmüller space of $\Sigma$ is trivial, actually $\operatorname{E}(S,\rho_{j})=\operatorname{Area}(\Sigma,\rho_{j})$ (see Definition 1.9). By Theorem 1.10, $\psi_{j}$ is both harmonic and weakly conformal, namely it is a branched minimal immersion. Hence,
(38) $\operatorname{Area}(\mathbf{D}_{S},(\psi_{j})^{*}g_{\mathrm{Eucl}})=\operatorname{E}(\psi_{j}|_{\mathbf{D}_{S}})=\operatorname{E}(S,\rho_{j})=\operatorname{Area}(\Sigma,\rho_{j}).$
Since $\rho_{j}(\pi_{1}(S))$ is finite, $\ker(\rho_{j})$ has finite index in $\pi_{1}(S)$. Thus, we can consider the connected finite covering $S_{j}$ of $S$, with fundamental group $\ker(\rho_{j})$. The map $\psi_{j}$ factors through a branched minimal immersion $\psi^{\prime}_{j}:S_{j}\to\mathbb{S}^{2N_{j}-1}$ with finite total energy. Let $\bar{S}_{j}$ be the closed Riemann surface obtained from $S_{j}$ by closing the punctures. By the removable singularity theorem of Sacks-Uhlenbeck [SU81] [CM11, Theorem 4.26], $\psi^{\prime}_{j}$ extends across the punctures to a branched minimal immersion still denoted by
$\psi^{\prime}_{j}:\bar{S}_{j}\to\mathbb{S}^{2N_{j}-1}.$
As a consequence,
$\psi_{j}(\tilde{S})=\psi^{\prime}_{j}(\bar{S}_{j})$
is a closed, branched, minimal surface in $\mathbb{S}^{2N_{j}-1}$.
By Theorem 0.4, the pullback metrics $(\psi_{j})^{*}g_{\mathrm{Eucl}}$, which descends to a metric on $S$ by equivariance of $\psi_{j}$, converges smoothly to $\frac{1}{8}g_{\mathrm{hyp}}$ on any compact subsets of $S$. Hence, the Gaussian curvature of $(\psi_{j})^{*}g_{\mathrm{Eucl}}$ converges smoothly to $-8$ away from the punctures. It remains to control the Gaussian curvature around the punctures, which is achieved with general properties of minimal surfaces. Let $\alpha_{j}$ be the degree of the covering $S_{j}\to S$. Let $\delta\in(0,1]$ be a small constant. Let ${A}_{\delta}\subset{S}$, $\tilde{A}_{\delta}\subset\tilde{S}$ be defined as in (14), and let $\tilde{A}_{j}$ be the image of $\tilde{A}_{\delta}$ in $S_{j}$ by the projection $\tilde{S}\to S_{j}$. Write
$\tilde{Z}_{j}:=\bar{S}_{j}\setminus\tilde{A}_{j}.$
Both $\tilde{A}_{j}$ and $\tilde{Z}_{j}$ implicitly depend on $\delta$. By (38), and Theorem 0.4,
(39) $\begin{split}&\lim_{j\to\infty}\frac{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}{\alpha_{j}}=\frac{\pi}{4},\\ &\lim_{\delta\to 0}\liminf_{j\to\infty}\frac{\operatorname{Area}(\tilde{A}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}=1,\\ &\lim_{\delta\to 0}\limsup_{j\to\infty}\frac{\operatorname{Area}(\tilde{Z}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}=0,\end{split}$
(40) $\lim_{j\to\infty}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{A}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}+8|=0.$
Note that $\partial A_{\delta}$ is a union of embedded curves around punctures with geodesic curvature constant equal to $8$ with respect to the rescaled hyperbolic metric $\frac{1}{8}g_{\mathrm{hyp}}$ on $S$. Thus by (36) again, as $j\to\infty$, the geodesic curvature $\kappa$ of $\partial\tilde{Z}_{j}$ inside of $(\tilde{Z}_{N},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})$ becomes arbitrarily close to the constant $8$. Besides,
$\lim_{j\to\infty}\frac{1}{\alpha_{j}}\operatorname{Length}(\partial\tilde{Z}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})=\lim_{j\to\infty}\frac{1}{\alpha_{j}}\operatorname{Length}(\partial\tilde{A}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})=\operatorname{Length}(\partial{A}_{\delta},\frac{1}{8}g_{\mathrm{hyp}})$
Thus, by (39) and since $\lim_{\delta\to 0}\operatorname{Length}(\partial{A}_{\delta},\frac{1}{8}g_{\mathrm{hyp}})=0$,
(41) $\liminf_{\delta\to 0}\limsup_{j\to\infty}\big{|}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\partial\tilde{Z}_{j}}\kappa\big{|}=0.$
The restriction of the map $\psi^{\prime}_{j}$ from $\tilde{Z}_{j}$ to $\mathbb{S}^{2N_{j}-1}$ is a branched minimal immersion of a disjoint union of disks. By the Gauss-Bonnet formula applied to $\tilde{Z}_{j}$,
$\int_{\tilde{Z}_{j}}K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}\geq-\int_{\partial\tilde{Z}_{j}}\kappa,$
where the inequality comes from the branched points and the disk topology (an easy way to see this inequality is to observe that each branched point can be locally smoothed in a neighborhood to a smooth immersion such that the integral of the Gaussian curvature in this neighborhood is negative). From (41), we get
(42) $\liminf_{\delta\to 0}\liminf_{j\to\infty}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}\geq 0.$
Next, by the Gauss equation [DCFF92, Chapter 6, Theorem 2.5], the Gaussian curvature (which is well-defined almost everywhere) satisfies
$K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}(x)\leq 1\quad\text{for almost all$x\in\bar{S}_{j}$}.$
This implies that
$\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}|$
$\leq\frac{\operatorname{Area}(\tilde{Z}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}+\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}(1-K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}})$
$=2\frac{\operatorname{Area}(\tilde{Z}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}-\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}.$
So combining (39) and (42, we find
$\limsup_{\delta\to 0}\limsup_{j\to\infty}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}|=0.$
Together with (39) and (40), we get
$0\leq$ $\limsup_{j\to\infty}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\bar{S}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}+8|$
$\leq$ $\limsup_{\delta\to 0}\limsup_{j\to\infty}\big{(}\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{A}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}+8|$
$+\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\tilde{Z}_{j}}(|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}|+8)\big{)}$
$=$ $0$
This almost ends the proof of (1): the branched minimal immersion $\psi^{\prime}_{j}:\bar{S}_{j}\to\mathbb{S}^{2N_{j}-1}$ satisfies all the requirements except that it is maybe not simple. But we can always find another branched minimal immersion $\varphi_{j}:T_{j}\to\mathbb{S}^{2N_{j}-1}$ where $T_{j}$ is a closed Riemann surface covered by $\bar{S}_{j}$, which is simple and which has same image as $\psi^{\prime}_{j}$. Since clearly
$\frac{1}{\operatorname{Area}(\bar{S}_{j},(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}})}\int_{\bar{S}_{j}}|K_{(\psi^{\prime}_{j})^{*}g_{\mathrm{Eucl}}}+8|=\frac{1}{\operatorname{Area}(T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}})}\int_{T_{j}}|K_{\varphi_{j}^{*}g_{\mathrm{Eucl}}}+8|,$
we conclude (1).
For (2): We continue with the notations used above. Recall that
$\psi_{j}(\tilde{S})=\psi_{j}^{\prime}(\bar{S}_{j})=\varphi_{j}(T_{j})$
where $\psi_{j}$, $\psi_{j}^{\prime}$ and $\varphi_{j}$ are branched minimal immersions and $\varphi_{j}$ is simple. From the proof of Proposition 4.4 and Theorem 4.6, we can assume that $\psi_{j}$ smoothly converges (after composing with unitary matrices) on compact subsets to a smooth, minimal immersion $\psi_{\infty}:\tilde{S}\to\mathbb{S}_{H}$ where $\mathbb{S}_{H}$ is the unit sphere of the smallest Hilbert space containing $\psi_{\infty}(\tilde{S})$, and $\psi_{\infty}$ is equivariant with respect to a representation $\rho^{\prime}:\pi_{1}(S)\to\operatorname{End}(H)$ such that $\rho^{\prime}\sim\lambda_{\pi_{1}(S)}$, and $\psi_{\infty}$ is equivariantly area-minimizing. By (35), $(\psi_{\infty})^{*}g_{H}=\frac{1}{8}g_{\mathrm{hyp}}$ where $g_{\mathrm{hyp}}$ is the hyperbolic metric on $\tilde{S}$ given by the lift of the hyperbolic metric on the thrice-punctured sphere $S$. In particular for each $\delta>0$, $\psi_{j}$ is an immersion on $\tilde{A}_{\delta}$ for large $j$, and $(\psi_{j})^{*}g_{\mathrm{Eucl}}$ converges smoothly to $\frac{1}{8}g_{\mathrm{hyp}}$ on $\tilde{A}_{\delta}$. By (39), the ratio of the $(\psi_{j})^{*}g_{\mathrm{Eucl}}$-area of $\tilde{A}_{j}$ and the $(\psi_{j})^{*}g_{\mathrm{Eucl}}$-area of $\bar{S}_{j}$ tends to $1$ as $j\to\infty$ and $\delta\to 0$.
We say that a branched immersion $\varphi$ from a domain $D$ is “almost injective” when it sends any two disjoint open sets $U,V\subset D$ to different images $\varphi(U)\neq\varphi(V)$. Due to the convergence properties of $(\psi_{j})^{*}g_{\mathrm{Eucl}}$ recalled above, if Benjamini-Schramm convergence of $(T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}})$ to $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$ fails, it must be because for some $R>0$, some $\delta>0$ and some $x_{j}\in\tilde{A}_{\delta}$, the restriction of $\psi_{j}$ to the $R$-ball in $(\tilde{S},(\psi_{j})^{*}g_{\mathrm{Eucl}})$ centered at $x_{j}$ is not almost injective for all $j$ large (i.e. $\psi_{j}$ maps the $R$-ball to a surface which intersects itself on some open set). Thus, since $\psi_{j}$ converges to $\psi_{\infty}$ on compact sets, in order to show Benjamini-Schramm convergence of $(T_{j},\varphi_{j}^{*}g_{\mathrm{Eucl}})$ to $(\mathbb{H}^{2},\frac{1}{8}g_{\mathrm{hyp}})$, we just need to show that the map $\psi_{\infty}$ is almost injective.
We now show that $\psi_{\infty}$ is almost injective. If that is not the case, then by the classical unique continuation property of minimal surfaces [CM11, Theorem 6.1] and the fact that the pullback metric $(\psi_{\infty})^{*}g_{H}$ is a rescaled hyperbolic metric, there is a nontrivial group of isometries $\mathcal{J}$ defined as the set of isometries $J$ of $(\tilde{S},\frac{1}{8}g_{\mathrm{hyp}})$ such that for any $x\subset\tilde{S}$, $\psi_{\infty}(J(x))=\psi_{\infty}(x)$. Since $\psi_{\infty}$ is an immersion, $\mathcal{J}$ acts freely on the rescaled hyperbolic plane $(\tilde{S},\frac{1}{8}g_{\mathrm{hyp}})$. Recall that $\pi_{1}(S)$ also acts by isometries on $(\tilde{S},\frac{1}{8}g_{\mathrm{hyp}})$. Let $\mathcal{J}^{\prime}$ be the subgroup of isometries generated by these two subgroups. Let us check that this group of isometries can neither be non-discrete, nor be discrete. Consider the surface $\tilde{S}/\mathcal{J}$ endowed with its rescaled hyperbolic metric $\frac{1}{8}g_{\mathrm{hyp}}$. This surface is not the hyperbolic plane because $\mathcal{J}$ is nontrivial. The group $\pi_{1}(S)$ is contained in the normalizer of $\mathcal{J}$ because if $J\in\mathcal{J}$ and $a\in\pi_{1}(S)$, then for any $x\subset\tilde{S}$,
$\psi_{\infty}(aJa^{-1}(x))=\rho^{\prime}(a)\big{(}\psi_{\infty}(Ja^{-1}(x))\big{)}=\rho^{\prime}(a)\big{(}\psi_{\infty}(a^{-1}(x))\big{)}=\psi_{\infty}(x).$
So $\pi_{1}(S)$ also acts by isometries on the rescaled hyperbolic surface $\tilde{S}/\mathcal{J}$. It is known that the isometry group of a complete (rescaled) hyperbolic surface, such as $\tilde{S}/\mathcal{J}$, is either discrete, or isomorphic to the circle $\mathbb{S}^{1}$ (in that case $\tilde{S}/\mathcal{J}$ is a topological annulus and $\mathcal{J}$ is abelian), see [ha]. In the latter case, the group $\pi_{1}(S)$ would have to act as a subgroup of the circle on the topological annulus $\tilde{S}/\mathcal{J}$, but this is impossible since $\pi_{1}(S)$ would have to be abelian. Hence $\pi_{1}(S)$ acts as a discrete subgroup of isometries on $\tilde{S}/\mathcal{J}$, and so $\mathcal{J}^{\prime}$ has to be discrete. But this is impossible too. Indeed if it is discrete, since $\pi_{1}(S)$ is a strict subgroup of $\mathcal{J}^{\prime}$, the quotient $\tilde{S}/\mathcal{J}^{\prime}$ is a nontrivial quotient of $(S,\frac{1}{8}g_{\mathrm{hyp}})$. But $S$ is the thrice-punctured sphere, which has no smooth nontrivial quotient, so $\tilde{S}/\mathcal{J}^{\prime}$ has at least one orbifold point with angle strictly less than $2\pi$. This means that for some nontrivial element $g\in\pi_{1}(S)$, the corresponding isometry of $\tilde{S}/\mathcal{J}$ is torsion. Since $H$ is by assumption the smallest Hilbert space containing $\psi_{\infty}(\tilde{S})$, the unitary transformation $\rho^{\prime}(g)$ of $H$ is uniquely determined by its action on $\psi_{\infty}(\tilde{S})$. We deduce that $\rho^{\prime}(g)$ is also torsion. But then, if $e\in\pi_{1}(S)$ is the identity element, for some positive integer $m$, $0=\|\rho^{\prime}(e)-\rho^{\prime}(g^{m})\|=\|\lambda_{\pi_{1}(S)}(e)-\lambda_{\pi_{1}(S)}(g^{m})\|\neq 0$ were the middle equality comes from Theorem 6.2 and the fact that $\rho^{\prime}\sim\lambda_{\pi_{1}(S)}$. This is a contradiction, thus $\mathcal{J}$ cannot be nontrivial, and $\psi_{\infty}$ is almost injective.
6. Appendix: About unitary representations
6.1. Weak containment and weak equivalence
All Hilbert spaces in this paper will be complex unless otherwise noted, and will have either finite or infinite countable dimensions. Let $\Gamma$ be a finitely generated group. A unitary representation $(\pi,H)$ of $\Gamma$ is a pair such that $H$ is a Hilbert space, and $\rho$ is a group morphism from $\Gamma$ to the unitary group $U(H)$ of $H$.
There is a natural topology on the set of all unitary representations of $\Gamma$, called the Fell topology [BdLHV08, Appendix F]. Convergence in the Fell topology can be described using the notion of weak containment [BdLHV08, Definition F.1.1]:
Definition 6.1.
Let $(\pi,H)$ and $(\rho,K)$ be two unitary representations of $\Gamma$. We say that $\pi$ is weakly contained in $\rho$ if for every $\xi\in H$, every finite subset $Q$ of $\Gamma$, and every $\epsilon>0$, there exist $\eta_{1},...,\eta_{m}$ in $K$ such that for all $g\in Q$,
$\big{|}\langle\pi(g)\xi,\xi\rangle-\sum_{j=1}^{m}\langle\rho(g)\eta_{j},\eta_{j}\rangle\big{|}<\epsilon.$
We write $\pi\prec\rho$ if the above holds. If we have both $\pi\prec\rho$ and $\rho\prec\pi$, then we say that $\pi$ and $\rho$ are weakly equivalent, and we write $\pi\sim\rho$.
Let $\mathbb{C}[\Gamma]$ be the set of complex, finite, linear combinations of elements of $\Gamma$. For any unitary representation of $\Gamma$ and $f\in\mathbb{C}[\Gamma]$, there is a well-defined continuous linear operator
$\pi(f):H\to H,$
see [BdLHV08, Section F.4]. Given such $\pi$ and $f$, let $\|\pi(f)\|$ denote the operator norm of $\pi(f)$.The next statement is contained in [BdLHV08, Theorem F.4.4]:
Theorem 6.2.
Let $(\pi,H)$ and $(\rho,K)$ be two unitary representations of $\Gamma$. The following are equivalent:
  1. (1)
    $\pi\prec\rho$
  2. (2)
    $\|\pi(f)\|\leq\|\rho(f)\|$ for all $f\in\mathbb{C}[\Gamma]$.
Let $\ell^{1}(\Gamma,\mathbb{C})$ be the set of $\mathbb{C}$-valued $\ell^{1}$-functions on $\Gamma$. An elementary fact is that $\|\pi(f)\|\leq\|f\|_{\ell^{1}}$ and so in the theorem above, “$f\in\mathbb{C}[\Gamma]$” can be replaced with “$f\in\ell^{1}(\Gamma,\mathbb{C})$”.
6.2. Strong convergence
Recall that the regular representation $\lambda_{\Gamma}:\Gamma\to\operatorname{End}(\ell^{2}(\Gamma,\mathbb{C}))$ of $\Gamma$ is the following canonical representation: for all $\gamma,x\in\Gamma$, $f\in\ell^{2}(\Gamma,\mathbb{C})$,
$(\lambda_{\Gamma}(\gamma).f)(x):=f(\gamma^{-1}x).$
Crucial to us is the notion of strong convergence:
Definition 6.3.
A sequence of unitary representations $\rho_{j}:\Gamma\to\operatorname{End}(V_{j})$ strongly converges to the regular representation if
$\forall z\in\mathbb{C}[\Gamma],\quad\lim_{j\to\infty}\|\rho_{j}(z)\|=\|\lambda_{\Gamma}(z)\|$
where $\lambda_{\Gamma}$ is the regular representation of $\Gamma$.
[…]
6.3. Examples of strongly convergent sequences
For the reader’s convenience, we list the results on strong convergence which are directly applied in this paper. In what follows, let $F_{k}$ be a free group of rank $k\geq 2$, generated by $x_{1},...,x_{k}$. Let $\lambda_{F_{k}}$ denote its (left) regular representation.
6.3.1. Random Haar unitaries
The first examples of unitary representations of free groups which strongly converge were constructed by Haagerup-Thorbjornsen [HT05]. These representations are defined by sending generators of the free group to random unitary matrices in $U(N)$ sampled with the probability measure which is the “exponential” of the GUE model [HT05, Theorem 8.2 and Remark 8.3]. The following result of Collins-Male [CM14, Theorem 1.5] confirms that the same conclusion holds when the sampling is done with the more obvious Haar measure (which is a different measure). For any $k$-tuple $(u_{1}^{(N)},...,u_{k}^{(N)})\in U(N)^{k}$, let $\rho_{(u_{1}^{(N)},...,u_{k}^{(N)})}:F_{k}\to U(N)$ be the unitary representation of $F_{k}$ where the generator $x_{i}$ is sent to $u_{i}^{(N)}$.
Theorem 6.4.
For each $N\geq 1$, let $u_{1}^{(N)},...,u_{k}^{(N)}$ be a family of independent random matrices in $U(N)$ sampled with the Haar measure. Then, almost surely, the sequence of unitary representations $\rho_{(u_{1}^{(N)},...,u_{k}^{(N)})}$ strongly converges.
6.3.2. Random permutations
For any integer $m\geq 3$, let $[m]:=\{1,...,m\}$ and $\mathbf{S}_{m}$ denote the permutation group of $[m]$. Let $V_{m}:=\ell^{2}([m],\mathbb{C})$ be the $m$-dimensional complex $\ell^{2}$-space endowed with the standard Euclidean metric, and $V^{0}_{m}\subset V_{m}$ the subspace of functions with average 00. The group $\mathbf{S}_{m}$ acts on $V_{m}$ by the standard unitary representation (via $0-1$ matrices):
[…]
and on $V^{0}_{m}$, which is the $(m-1)$-dimensional irreducible component, orthogonal to the diagonal direction:
[…]
For any $k$-tuple $(s_{1}^{(m)},...,s_{k}^{(m)})\in\mathbf{S}_{m}$, let $\rho_{(s_{1}^{(m)},...,s_{k}^{(m)})}:F_{k}\to\operatorname{End}(V^{0}_{m})$ be the unitary representation of $F_{k}$ equal to $\mathbf{std}\circ\theta_{(s_{1}^{(m)},...,s_{k}^{(m)})}$ where $\theta_{(s_{1}^{(m)},...,s_{k}^{(m)})}:F_{k}\to\mathbf{S}_{m}$ is the homomorphism that sends the generator $x_{i}$ to $s_{i}^{(m)}$. Below is a special case of the strong convergence theorem of Bordenave-Collins [BC19, Theorem 3]:
Theorem 6.5.
For each $m>1$, let $s_{1}^{(m)},...,s_{k}^{(m)}$ be a family of independent random permutations in $\mathbf{S}_{m}$ with respect to the uniform measure. Then, the sequence of unitary representations $\rho_{(s_{1}^{(m)},...,s_{k}^{(m)})}$ strongly converges in probability, in the following sense: for any $\varepsilon>0$, for any $z\in\mathbb{C}[\Gamma]$, with probability tending to $1$ as $m\to\infty$, we have
$\|\rho_{(s_{1}^{(m)},...,s_{k}^{(m)})}(z)\|\leq\|\lambda_{F_{k}}(z)\|+\varepsilon.$
We will also need the following elementary lemma, just in the special case $k=2$ and $F_{2}=\langle a,b\rangle$. Let $\mathbb{S}_{V^{0}_{m}}$ be the unit sphere of $V^{0}_{m}$ endowed with the standard Euclidean metric.
Lemma 6.6.
For each $m>1$, let $s_{1}^{(m)},s_{2}^{(m)}$ be two independent random permutations in $\mathbf{S}_{m}$ with respect to the uniform measure. With probability tending to $1$ as $m\to\infty$, the unitary representation $\rho_{m}:=\rho_{(s_{1}^{(m)},s_{2}^{(m)})}:F_{2}\to\operatorname{End}(V^{0}_{m})$ is such that $\rho_{m}(a)$, $\rho_{m}(b)$ and $\rho_{m}(b^{-1}a)$ all have a respective fixed point in $\mathbb{S}_{V^{0}_{m}}$.
Proof.
We claim that the unitary transformation $\rho_{m}(a)$ has a fixed point in $\mathbb{S}_{V^{0}_{m}}$ if and only if the corresponding permutation $s_{1}^{(m)}$ admits a nontrivial cycle decomposition, in the sense that it admits at least two cycles. Indeed, if $x,y\in[m]$ belong to two different cycles of $s_{1}^{(m)}$, then the subsets
$X:=\{(s_{1}^{(m)})^{l}(x)\}_{l\in\mathbb{Z}},\quad Y=\{(s_{1}^{(m)})^{l}(y)\}_{l\in\mathbb{Z}}$
are disjoint, and if $\sharp X,\sharp Y$ denote their sizes, one can define the function
$f(z)=\left\{\begin{array}[]{rcl}\sqrt{\frac{\sharp Y}{\sharp X.\sharp Y+(\sharp X)^{2}}}&\mbox{for}&z\in X\\ -\frac{\sharp X}{\sharp Y}\sqrt{\frac{\sharp Y}{\sharp X.\sharp Y+(\sharp X)^{2}}}&\mbox{for}&z\in Y\\ 0&\mbox{for}&z\in[m]\setminus(X\cup Y)\end{array}\right..$
One checks that $f$ has average 00 and $\ell^{2}$-norm $1$, so it is an element of $\mathbb{S}_{V^{0}_{m}}$. Moreover, it is indeed a fixed point of $\rho_{m}(a):=\mathbf{std}(s_{1}^{(m)})$. Conversely, if $s_{1}^{(m)}$ is made of one cycle, then it does not fix any nontrivial function with average 00 (via the standard representation). This shows the claim.
Next, since $s_{1}^{(m)}$ is chosen uniformly at random in $\mathbf{S}_{m}$, the probability that it admits a nontrivial cycle decomposition converges to $1$ as $m\to\infty$ (by a simple counting argument, the probability of a random permutation of $\mathbf{S}_{m}$ to have a cycle of length $m$ is $\frac{1}{m}$). Next, $s_{2}^{(m)}$ is chosen independently of $s_{1}^{(m)}$ and so $(s_{2}^{(m)})^{-1}s_{1}^{(m)}$ is also a uniformly random element of $\mathbf{S}_{m}$. Thus, $\theta_{m}(a)$, $\theta_{m}(b)$, $\theta_{m}(b^{-1}a)$ all have a nontrivial cycle decomposition with probability tending to $1$. We conclude with the claim above.
6.3.3. Representations of surface groups
Based on the paper of Bordenave-Collins [BC19], Louder-Magee [LMH25, Corollary 1.2] constructed strongly convergent sequence of unitary representations for limit groups, in particular surface groups.
Theorem 6.7.
Let $\Gamma$ be the fundamental group of a closed oriented surface. There exists a sequence of unitary representations $\rho_{j}:\Gamma\to U(N_{j})$ which strongly converges, and such that $\rho_{j}(\Gamma)$ is a finite subgroup of $U(N_{j})$ for any $j\geq 1$.